### v1

# Relatively big projective modules and their applications to direct sum decompositions

Abstract.
Countably generated projective modules that are relatively big with respect to a trace ideal were introduced by P. Příhoda, as an extension of Bass’ uniformly big projectives. It has already been proved that there are a number of interesting examples of rings whose countably generated projective modules are always relatively big. In this paper, we increase the list of such examples, showing that it includes all right noetherian rings satisfying a polynomial identity.
We also show that countably generated projective modules over locally semiperfect torsion-free algebras over $h$-local domains are always relatively big. This last result applies to endomorphism rings of finitely generated torsion-free modules over $h$-local domains. As a consequence, we can give a complete characterization of those $h$-local domains of Krull dimension $1$ for which every direct summand of a direct sum of copies of a single finitely generated torsion-free module is again a direct sum of finitely generated modules.
The first and second authors were partially supported by the projects MIMECO PID2023-147110NB-I00 financed by the Spanish Government and Laboratori d’Interaccions entre Geometria, Àlgebra i Topologia (LIGAT) with reference number 2021 SGR 01015 financed by the Generalitat de Catalunya.
The third author was supported by Czech Science Foundation grant GAČR 23-05148S
Introduction
For a long time, the study of projective modules was focused on the finitely generated ones. Outside the finitely generated setting, the classical theorem of Kaplansky [kaplansky] stating that every projective module is a direct sum of countably generated projective modules shows that one can just consider the countably generated case. Interesting monoids associated to infinitely generated projective modules have recently been studied in several different contexts: the monoid $V(\Lambda)\sqcup B(\Lambda)$ of relatively big projectives defined, formally, in [wiegand] and that is the main object of interest in this paper, but also the analogous of the Cuntz monoid for general rings defined in [AABPV], and the $\kappa$-monoids defined in [nazemian_smertnig].
Let $\Lambda$ be a ring, and let $V(\Lambda)$ (resp. $V^{*}(\Lambda)$) denote the set of representatives of the isomorphism classes of finitely generated (resp. countably generated) projective right $\Lambda$-modules. These sets form commutative monoids with addition induced by the direct sum of modules, with $V(\Lambda)$ naturally embedded in $V^{*}(\Lambda)$.
The concept of relatively big projective module was introduced by Příhoda in [P2], extending Bass’ uniformly big projectives [bass]. A projective right $\Lambda$-module $P$ is said to be $J$-big, for $J$ an ideal of $\Lambda$, if every countably generated projective $\Lambda$-module whose trace is contained in $J$ is a homomorphic image of $P$. Every $J$-big projective module $P$ is determined by the ideal $J$ and the isomorphism class of $P/PJ$ (cf. Lemma 1.14). A countably generated projective module $P$ is said to be relatively big if there exist $J$ such that $P$ is $J$-big and $P/PJ$ is finitely generated.
To capture the properties of relatively big projectives, we introduced in [wiegand] the monoid $V(\Lambda)\sqcup B(\Lambda)$ of isomorphism classes of relatively big projective modules. This monoid embeds in $V^{*}(\Lambda)$ and can be thought of as the disjoint union of $V(\Lambda/I)$, where $I$ runs through the trace ideals of the countably generated $\Lambda$-modules, see Corollary 2.7.
In [P2], it was proved that, over a noetherian ring $\Lambda$ which satisfies a certain descending condition on two-sided ideals, that we call the ()(*)( ∗ )-condition, $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ (cf. [P2, Lemma 2.8] and Theorem 1.18). Commutative noetherian rings, finitely generated algebras over commutative noetherian rings of Krull dimension one and noetherian semilocal rings satisfy the ()(*)( ∗ ) condition, therefore all countably generated projective modules over these rings are relatively big. These results were the main ingredient to make a detailed study of projective modules over noetherian semilocal rings in [crelle] or to study some infinite direct sum decompositions in [wiegand].
The main results of this paper give new classes of rings for which $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$. We prove that right noetherian PI-rings satisfy the ()(*)( ∗ )-condition (see Theorem 4.8), and this is one of the main ingredients to show all countably generated right (or left) projective modules over a right noetherian PI-ring are relatively big in Corollary 4.9. In particular, this applies to module-finite algebras over commutative noetherian rings, such as group algebras of finite groups over commutative noetherian rings or to endomorphism rings of finitely generated modules over commutative noetherian rings, so we show that the restriction on Krull dimension we were imposing in our previous works was superfluous.
Outside the noetherian setting, it is easy to show that for semiperfect rings all countably generated projective right (or left) modules are relatively big (cf. Proposition 6.3). In Theorem 6.16, we show that the same conclusion holds for $\Lambda$ a locally semiperfect torsion-free algebra, over an $h$-local domain $R$ with field of fractions $Q$, such that $\Lambda\otimes_{R}Q$ is simple artinian.
If $\Lambda$ is the endomorphism ring of a finitely generated torsion-free module $M$ over an $h$-local domain $R$ then $\Lambda\otimes_{R}Q$ is simple artinian. If, in addition, $R$ has Krull dimension $1$ and all the direct summands of a direct sum of copies of $R\oplus M$ are again a direct sum of finitely generated modules, it was proved in [AHP] that $\Lambda$ is locally semiperfect. The converse of this result is not true, in general, as additional conditions in the ranks of indecomposable modules over the localizations of $R$ at its maximal ideals must be satisfied. Our results on projective modules allow us to prove that such rank conditions are not only necessary but also sufficient to prove the converse (cf. Theorem 9.3). This result complements and clarifies the results in [AHP].
It is interesting to recall that Procesi and Small [PS] proved that the endomorphism ring of a finitely generated module over a commutative ring or, in general, over a ring satisfying a polynomial identity, is also a ring satisfying a polynomial identity. As far as we know, this point of view has not been so much examined in the world of finitely generated modules, endomorphism rings and direct sum decompositions and, certainly, it seems it should be worthwhile exploring it.
The paper is structured as follows. In Section 1 we give the basic definitions and notation. We introduce the needed background, presenting the concept of relatively big projective, the ()(*)( ∗ )-condition and the main results relating both notions. Section 2 we examine the monoid of isomorphism classes of relatively big projective modules, its algebraic properties and explore the right/left (non)symmetry of the monoids; check Example 2.10 and Corollary 2.12. In Section 3 we show that countably generated, projective right modules lift modulo a right $T$-nilpotent ideal, cf. Lemma 3.1. We prove that for a right $T$-nilpotent ideal $N$ of $\Lambda$ there are monoid isomorphisms between $V^{*}(\Lambda)$ and $V^{*}(\Lambda/N)$. This isomorphism restricts to isomorphisms between $V(\Lambda)$ and $V(\Lambda/N)$, and between $V(\Lambda)\sqcup B(\Lambda)$ and $V(\Lambda/N)\sqcup B(\Lambda/N)$, cf. Corollary 3.3. In Section 4, these results are an important ingredient to prove that right noetherian PI-rings satisfy the ()(*)( ∗ )-condition (Theorem 4.8), and to conclude that $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ for right noetherian PI-rings in Corollary 4.9. In particular, this holds for module-finite algebras over commutative noetherian rings (cf. Corollary 4.10).
In the rest of the paper, we turn to the study of algebras over $h$-local domains. These are domains satisfying strong noetherianity conditions on the maximal spectrum, more precisely, any non-zero ideal of $R$ is contained only in a finite number of maximal ideals, and any non-zero prime ideal of $R$ is contained in a unique maximal ideal. In particular, noetherian domains of Krull dimension one are always $h$-local. The definition of $h$-local domain is attributed to Matlis. However, Jaffard in [jaffard] introduced Dedekind-type domains as those domains $R$ such that any nonzero ideal is a finite product of ideals that are contained in a unique maximal ideal of $R$. Jaffard proved that Dedekind-type domains coincide with $h$-local domains, cf. [FHL, Theorem 2.1.5].
The class of $h$-local domains is ubiquitously associated with many classical problems in the study of modules over non-noetherian domains, see [matlis3] and the recent survey [olberding]. Non-finitely generated projective modules over an $h$-local domain are free [hinohara], so what is of interest is the description of projective modules on algebras over such domains.
Our main motivation to study this setting comes from our paper [AHP], in which we studied domains for which the class of direct sums of finitely generated torsion-free modules is closed under direct summands. The type of algebras over $h$-local domains we consider appears naturally when considering this question. We will find some neat applications to this question of the theory we develop in the final section of the paper.
Section 5 discusses Package Deal Theorems for projective modules over $\Lambda$, whenever $\Lambda$ is an algebra over an $h$-local domain; our goal is twofold, recalling the results from [AHP] and proving some extension needed in the subsequent sections. Section 6 focuses on semiperfect and locally semiperfect algebras. We observe in Proposition 6.3 that $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ for semiperfect rings.
Let $R$ be an $h$ local domain with field of fractions $Q$. In the final part of the paper, we study locally semiperfect algebras $\Lambda$ over $R$ such that $\Lambda\otimes_{R}Q$ is simple artinian. In Section 6, we show that every trace of a projective module is the trace of a finitely generated projective module (cf. Lemma 6.6) and $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ (cf. Theorem 6.16). In Section 7, we study in detail the genus of projective modules over such algebras that we understand as a submonoid of the product of $V^{*}(\Lambda_{\mathfrak{m}})$ for $\mathfrak{m}$ maximal ideal of $R$ that can be described quite explicitly. In Proposition 7.1, we show that two countably generated projective modules that are not finitely generated are isomorphic if and only if all their localizations at maximal ideals of $R$ are isomorphic. That is, they are isomorphic if and only if they have the same genus.
These tools allow us to study in Section 8 when every projective module is a direct sum of finitely generated ones. Finally, in Section 9, we apply our results to $\Lambda$ the endomorphism ring of a finitely generated torsion-free module $M$ over an $h$-local domain $R$ of Krull dimension $1$. In Theorem 9.3, we determine when any direct summand of an arbitrary direct sum of copies of $M$ is a sum of finitely generated modules. The results in this section complement the ones obtained in [AHP] in which we were studying domains for which the class of direct sums of finitely generated torsion-free modules is closed under direct summands.
Throughout the paper, rings are associative with $1$, and morphisms are unital. In each section, we try to be very precise about the setting we are working with. We mainly use $R$ to denote a commutative ring and $\Lambda$ to denote a non-necessarily commutative ring or an algebra over a domain $R$.
1. The monoid of infinitely generated projective modules
In this section we will cover the basic notions on projective modules and their trace ideals. We start by introducing the notation already used in [wiegand] for monoids of projective modules.
Notation 1.1.
For a given ring $\Lambda$, we write $V_{r}(\Lambda)$ (resp. $V_{r}^{*}(\Lambda)$) for the set of representatives of the isomorphism classes of finitely generated (resp. countably generated) projective right $\Lambda$-modules. If $P$ is a projective right $\Lambda$-module in $\mbox{\rm add}(\Lambda)$ (resp. in $\mbox{\rm Add}(\Lambda)$), we denote its representative in $V_{r}(\Lambda)$ (resp. in $V_{r}^{*}(\Lambda)$) by $[P]$.
Similarly, we write $V_{\ell}(\Lambda)$ (resp. $V_{\ell}^{*}(\Lambda)$) for the corresponding sets of representatives of the isomorphism classes of projective left $\Lambda$-modules and follow the same notation for the representatives of the isomorphism classes in these sets.
Unless stated otherwise, our modules will always be right modules, so we will write $V(\Lambda)$ and $V^{*}(\Lambda)$ instead of $V_{r}(\Lambda)$ and $V_{r}^{*}(\Lambda)$, respectively, and leave the previous notation in cases where both right and left modules appear. Every result stated for $V(\Lambda)$ and $V^{*}(\Lambda)$ can be stated in the context of left $\Lambda$-modules for $V_{\ell}(\Lambda)$ and $V_{\ell}^{*}(\Lambda)$, respectively.
Remark 1.2.
Note that both $V(\Lambda)$ and $V^{*}(\Lambda)$ are commutative monoids with addition defined by $[P]+[Q]=[P\oplus Q]$. Note that $V(\Lambda)$ is a submonoid of $V^{*}(\Lambda)$ and that $V^{*}(\Lambda)\setminus V(\Lambda)$ is a subsemigroup of $V^{*}(\Lambda)$.
An important object related to projective modules is trace ideals. We recall the definition and some basic properties.
Definition 1.3.
Let $\Lambda$ be a ring, and let $M$ be a right $\Lambda$-module. Then the trace of $M$ in $\Lambda$ is the two-sided ideal of $\Lambda$ defined as
$\operatorname{Tr}_{\Lambda}(M)=\sum_{f\in\operatorname{Hom}_{\Lambda}(M,\Lambda)}f(M).$
Alternatively, if $S=\mathrm{End}_{\Lambda}(M)$ the trace of $M$ is the image of the pairing
We will denote by $\mathcal{T}_{r}(\Lambda)$ (resp. $\mathcal{T}_{\ell}(\Lambda)$) the set of ideals of $\Lambda$ that are traces of countably generated projective right (resp. left) $\Lambda$-modules.
Remark 1.4.
Let $\Lambda$ be a ring and let $M$ be a right $\Lambda$-module with endomorphism ring $S=\operatorname{End}_{\Lambda}(M)$. Then $M$ is also a left $S$-module. Since $\operatorname{Hom}_{\Lambda}(M,\Lambda)$ is both a right $S$-module and a left $\Lambda$-module, the tensor product $\operatorname{Hom}_{\Lambda}(M,\Lambda)\otimes_{S}M$ is well-defined. Then $\operatorname{Tr}_{\Lambda}(M)$ is the image of the trace map:
${\tau\colon}$${\operatorname{Hom}_{\Lambda}(M,\Lambda)\otimes_{S}M}$${\Lambda}$${f\otimes x}$${f(x).}$
In the following, we will write $\mathcal{T}(\Lambda)$ instead of $\mathcal{T}_{r}(\Lambda)$, and leave the previous notation in cases where both right and left modules appear.
Lemma 1.5.
Let $\Lambda$ be a ring, let $M,N$ be right $\Lambda$-modules, let $P$ be a projective right $\Lambda$-module, and let $J$ be an ideal of $\Lambda$. Then
  1. (i)
    If $M\subseteq_{\oplus}N$, then $\operatorname{Tr}_{\Lambda}(M)\subseteq\operatorname{Tr}_{\Lambda}(N)$.
  2. (ii)
    $\operatorname{Tr}_{\Lambda}(M\oplus N)=\operatorname{Tr}_{\Lambda}(M)+\operatorname{Tr}_{\Lambda}(N)$.
  3. (iii)
    [lam2, Proposition 2.40] $\operatorname{Tr}_{\Lambda}(P)\subseteq J$ if and only if $PJ=P$. In addition, the trace of a projective module is always an idempotent ideal.
We recall the following properties of traces of projective modules.
Lemma 1.6.
Let $\Lambda$ be a ring, let $P$ be a right $\Lambda$-module, and let $J$ be an ideal of $\Lambda$. Then $(\operatorname{Tr}_{\Lambda}(P)+J)/J\subseteq\operatorname{Tr}_{\Lambda/J}(P/PJ)$. Moreover, if $P$ is projective, then $(\operatorname{Tr}_{\Lambda}(P)+J)/J=\operatorname{Tr}_{\Lambda/J}(P/PJ)$.
Remark 1.7.
For a commutative ring $\Lambda$, Vasconcelos showed that trace ideals of projective modules are exactly the pure ideals.
In [whitehead], Whitehead gave a characterization of ideals that are traces of projectives for general rings. In particular, he showed that an idempotent ideal of $\Lambda$ that is finitely generated as a left ideal is always the trace of a countably generated projective right $\Lambda$-module. We shall use this result throughout the paper, sometimes without prior acknowledgment.
An alternative approach to these results characterizing the traces of projective modules can also be found in the paper [traces].
Lemma 1.8.
Let $\Lambda$ be a ring with Jacobson radical $J(\Lambda)$. Let $I$ and $I^{\prime}$ be elements in $\mathcal{T}(\Lambda)$. Then $I=I^{\prime}$ if and only if $I+J(\Lambda)=I^{\prime}+J(\Lambda)$.
Therefore, if $I\in\mathcal{T}(\Lambda)$ then $I+J(\Lambda)/J(\Lambda)\in\mathcal{T}(\Lambda/J(\Lambda))$, and the assignment defined by $I\mapsto(I+J(\Lambda))/J(\Lambda)$ defines an injective map $\varphi\colon\mathcal{T}(\Lambda)\to\mathcal{T}(\Lambda/J(\Lambda))$.
Proof.  The first part of the statement is included in [traces, Corollary 2.9].
To prove the second part of the statement, note that if $I$ is the trace of the projective right $\Lambda$-module $P$ then, by Lemma 1.6, $I+J(\Lambda)/J(\Lambda)$ is the trace of $P/PJ(\Lambda)$. This shows that $\varphi$ is well defined. The injectivity of $\varphi$ follows from the first part of the statement.
The following lemma states that projective modules can be lifted modulo a trace ideal.
Lemma 1.9.
[traces, Theorem 3.1, Corollary 3.2] Let $\Lambda$ be a ring, and let $I$ be the trace of a projective right $\Lambda$-module. Let $P^{\prime}$ be a projective right $\Lambda/I$-module. Then there is a projective right $\Lambda$-module $X$ such that $I\subseteq\operatorname{Tr}_{\Lambda}(X)$, $X/XI\cong P^{\prime}$ and $\operatorname{Tr}_{\Lambda}(X)/I=\operatorname{Tr}_{\Lambda/I}(P^{\prime})$.
Moreover, if $I$ is an element in $\mathcal{T}(\Lambda)$ and $P^{\prime}$ is countably generated, then $X$ can be taken to be countably generated.
1.1. Relatively big projectives
We introduce the notion of relatively big projective modules with respect to an ideal. This definition was first introduced by P. Příhoda in [P2] as a way of generalizing Bass’ uniformly big projective modules. We motivate the definition of this class of projective modules by comparing it with a characterization of the class of modules generated by a projective module.
Definition 1.10.
Let $\Lambda$ be a ring, and let $P$ be a projective right $\Lambda$-module. We define the class $\mathrm{Gen}(P)$ of modules generated by $P$ as the class of right $\Lambda$-modules which are a homomorphic image of a direct sum of copies of $P$.
Lemma 1.11.
[wiegand, Lemma 2.10] Let $\Lambda$ be a ring, and let $P$ be a projective right $\Lambda$-module with trace ideal $I$. Then
$\mathrm{Gen}(P)=\{M\in\mbox{\rm Mod-}\Lambda:MI=M\}.$
Definition 1.12.
Let $\Lambda$ be a ring, let $P$ be a countably generated projective right $\Lambda$-module, and let $J$ be an ideal of $\Lambda$. The module $P$ is said to be big with respect to $J$, or $J$-big, if for every countably generated projective right $\Lambda$-module $Q$ satisfying $QJ=Q$, then $Q$ is a homomorphic image of $P$.
If $P$ is $J$-big projective and $J$ does not contain a nonzero element of $\mathcal{T}(\Lambda)$, then Definition 1.12 is only fulfilled for the trivial case $Q=\{0\}$. If $0\neq K\in\mathcal{T}(\Lambda)$ is such that $K\subseteq J$, then $P$ is $K$-big.
We summarize what it means to be relatively big with respect to an ideal which is the trace of a countably generated projective module.
Lemma 1.13.
Let $\Lambda$ be a ring, and let $P,Q$ be countably generated projective right $\Lambda$-modules with trace ideals $I$ and $J$, respectively. Then
  1. (i)
    $P$ is $J$-big if and only if $P\oplus Q^{(\omega)}\cong P$.
  2. (ii)
    If $P$ is $J$-big such that $P/PJ$ is finitely generated, then $J$ is the minimal ideal of $\Lambda$ such that $P/PJ$ is finitely generated.
Proof.  $(i)$. Since $J=\operatorname{Tr}_{\Lambda}(Q)=\operatorname{Tr}_{\Lambda}(Q^{(\omega)})$, by Lemma 1.5, $Q^{(\omega)}J=Q^{(\omega)}$. If $P$ is $J$-big, then there is an epimorphism $P\to Q^{(\omega)}$. Since $Q^{(\omega)}$ is projective, this epimorphism splits, so $Q^{(\omega)}$ is isomorphic to a direct summand of $P$. Write $P\cong P^{\prime}\oplus Q^{(\omega)}$. Then, $P\cong P^{\prime}\oplus Q^{(\omega)}\cong P^{\prime}\oplus Q^{(\omega)}\oplus Q^{(\omega)}\cong P\oplus Q^{(\omega)}$.
Conversely, if $P\oplus Q^{(\omega)}\cong P$, then $Q^{(\omega)}$ is isomorphic to a direct summand of $P$, so there exists an epimorphism $P\to Q^{(\omega)}$. By Lemma 1.5, $QJ=Q$, so $P$ is $J$-big.
Statement $(ii)$ is [wiegand, Lemma 3.3].
Lemma 1.14.
[wiegand, Proposition 3.4] Let $\Lambda$ be a ring, let $J\in\mathcal{T}(\Lambda)$, and let $P$ be a $J$-big projective right $\Lambda$-module. Then the isomorphism class of $P$ is determined by the ideal $J$ and the isomorphism class of $P/PJ$, i.e., if $P^{\prime}$ is a $J$-big projective right $\Lambda$-module and $P^{\prime}/P^{\prime}J\cong P/PJ$ then $P^{\prime}\cong P$.
1.2. Finite generation modulo an ideal and the ()(*)( ∗ )-condition
Let $\Lambda$ be a ring, and let $M$ be a right $\Lambda$-module. Consider the set
$\mathcal{I}(M)=\{I\unlhd\Lambda\mid M/MI\mbox{ is finitely generated}\}$
We recall the following crucial properties of a descending sequence of right ideals associated to a column finite idempotent matrix. The result is taken from [traces], but all the needed ideas are already in [P2].
Lemma 1.15.
[traces, Lemmas 5.1 and 5.2] Let $\Lambda$ be a ring. Let $A=(a_{i,j})$ be a countable column finite matrix with entries in $\Lambda$ such that $A^{2}=A$, and set $P=A\Lambda^{(\mathbb{N})}$. For any $k\geq 0$, let $I_{k}=\sum_{i>k,j\in\mathbb{N}}a_{ij}\Lambda$. Then,
  1. (i)
    $\operatorname{Tr}_{\Lambda}(P)=\Lambda I_{0}$.
  2. (ii)
    For any $k\geq 0$ and any $0<i\leq k$,
    $\sum_{j\in\mathbb{N}}a_{ij}\Lambda+\Lambda I_{k}/\Lambda I_{k}=\sum_{1\leq\ell\leq k}a_{i\ell}\Lambda+\Lambda I_{k}/\Lambda I_{k}.$
    In particular, if $0\leq n\leq k$, $I_{n}+\Lambda I_{k}/\Lambda I_{k}$ is finitely generated.
  3. (iii)
    For any $k\geq 0$, there exists $n_{k}>k$ such that $I_{n_{k}}I_{k}=I_{n_{k}}$. In particular, the descending sequence of two-sided ideals $(\Lambda I_{k})_{k\geq 0}$ has a subsequence $(J_{n})_{n\geq 0}$ of two-sided ideals satisfying that:
    • (a)
      for any $n\geq 0$, $J_{n+1}J_{n}=J_{n+1}$;
    • (b)
      for any $k\geq 0$, there exists $n$ and $\ell$ such that $J_{n}=\Lambda I_{k+\ell}\subseteq\Lambda I_{k}$.
  4. (iv)
    For any $k\geq 0$, $\Lambda I_{k}\in\mathcal{I}(P)$. Moreover, for any $I\in\mathcal{I}(P)$, there exists $k_{0}$ such that $\Lambda I_{k_{0}}\leq I$.
  5. (v)
    $\mathcal{I}(P)$ has minimal elements if and only if there exists $k_{0}\in\mathbb{N}$ such that, for any $\ell\in\mathbb{N}$, $\Lambda I_{k_{0}}=\Lambda I_{k_{0}+\ell}$. In this case $I_{k_{0}}=I_{k_{0}}^{2}$ and $\Lambda I_{k_{0}}$ is the minimal element of $\mathcal{I}(P)$.
Statement $(v)$ in Lemma 1.15 motivates the following definition.
Definition 1.16.
[P2] A ring $\Lambda$ satisfies condition ()(*)( ∗ ) provided every descending chain of two-sided ideals
$J_{1}\supseteq J_{2}\supseteq\cdots\supseteq J_{k}\supseteq\cdots$
such that $J_{k+1}J_{k}=J_{k+1}$, for each $k\geq 1$, is eventually stationary.
The following statement explains the role of condition ()(*)( ∗ ) in our discussion.
Corollary 1.17.
Let $\Lambda$ be a ring. Let $P$ be a countably generated projective right $\Lambda$-module. If $\Lambda$ satisfies condition ()(*)( ∗ ), then $\mathcal{I}(P)$ has a minimal element and this minimal element is an idempotent ideal.
Proof.  By Lemma 1.15 $(v)$, $\mathcal{I}(P)$ has a minimal element if and only if the descending sequence of two-sided ideals $(\Lambda I_{k})_{k\geq 0}$ is stationary. By Lemma 1.15 $(iii)$, this will happen if and only if the partial sub-sequence $(J_{n})_{n\geq 0}$ given by that statement is stationary. Since such partial sub-sequence satisfies that $J_{n+1}J_{n}=J_{n+1}$, for any $n\geq 0$, it is stationary because $\Lambda$ satisfies condition ()(*)( ∗ ).
The fact that the minimal element in $\mathcal{I}(P)$ is an idempotent ideal follows from Lemma 1.15 $(v)$.
Theorem 1.18.
[P2, Lemma 2.8] Let $\Lambda$ be a noetherian ring. Let $P$ be a countably generated projective right $\Lambda$-module such that $\mathcal{I}(P)$ has a minimal element $J$. Then $J$ is the trace ideal of a countably generated projective right $\Lambda$-module, and $P$ is $J$-big. In particular, if $\Lambda$ satisfies condition ()(*)( ∗ ), then for any countably generated projective right module $P$, there exists an ideal $J$ such that:
  • (i)
    $J$ is the trace of a countably generated projective right $\Lambda$-module,
  • (ii)
    $P$ is $J$-big, and
  • (iii)
    $P/PJ$ is finitely generated.
Remarks 1.19.
$(1)$ Let $\Lambda$ be a semilocal ring, and let $P$ be a countably generated projective right $\Lambda$-module. Then, by [traces, Proposition 5.6], $\mathcal{I}(P)$ has a minimal element $J$. Then $J$ is an idempotent ideal, but it does not need to be the trace of a projective module as it can even be contained in the Jacobson radical.
$(2)$ We do not know whether condition ()(*)( ∗ ) is left-right symmetric, even for the case of noetherian rings.
$(3)$ There are noetherian rings with infinite descending sequences of idempotent ideals, so they do not satisfy condition ()(*)( ∗ ). Over such rings one constructs a countably generated projective module $P$ such that $\mathcal{I}(P)$ does not have a minimal element. See, for example, the last section of [traces].
2. The monoid of relatively big projective modules
We want to focus on $J$-big projective modules $P$ such that $P/PJ$ is finitely generated. This motivates the following definition.
Definition 2.1.
Let $\Lambda$ be a ring. We denote by $V(\Lambda)\sqcup B(\Lambda)$ the subset of $V^{*}(\Lambda)$ of isomorphism classes of countably generated projective right $\Lambda$-modules $P$ such that there exists a countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ satisfying:
  1. (i)
    $P$ is $J$-big, that is, $P\oplus Q^{(\omega)}\cong P$.
  2. (ii)
    $P/PJ$ is a finitely generated projective right $\Lambda/J$-module.
Remark 2.2.
Let $\Lambda$ be a ring. Then, $B(\Lambda)$ is the subset of $V^{*}(\Lambda)\setminus V(\Lambda)$ of isomorphism classes of countably generated projective right $\Lambda$-modules $P$, which are not finitely generated, and such that there exists a non-zero countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ satisfying $(i)$ and $(ii)$ in the previous definition.
Note that a countably generated projective $\Lambda$-module whose isomorphism class is in $B(\Lambda)$ is never indecomposable.
Example 2.3.
Let $\Lambda$ be a ring, let $P$ be a non-zero countably generated projective right $\Lambda$-module, and let $X$ be a finitely generated projective right $\Lambda$-module. Then
  1. (i)
    $[X]\in V(\Lambda)\sqcup B(\Lambda)$ with $Q=\{0\}$ and $J=\{0\}$.
  2. (ii)
    $[X\oplus P^{(\omega)}]\in B(\Lambda)$ with $Q=P$ and $J=\operatorname{Tr}_{\Lambda}(P)$.
Lemma 2.4.
Let $\Lambda$ be a ring. Then
  1. (i)
    $B(\Lambda)$ is a subsemigroup of $V^{*}(\Lambda)\setminus V(\Lambda)$.
  2. (ii)
    $V(\Lambda)\sqcup B(\Lambda)$ is a submonoid of $V^{*}(\Lambda)$.
Proof.  $(i)$. Let $[P_{1}],[P_{2}]\in B(\Lambda)$. Then, for each $i=1,2$, there exists a countably generated projective right $\Lambda$-module $Q_{i}$ with trace ideal $J_{i}$ such that $P_{i}\oplus Q_{i}^{(\omega)}\cong P_{i}$ and $P_{i}/P_{i}J_{i}$ is a finitely generated projective right $\Lambda/J_{i}$-module. Take $P=P_{1}\oplus P_{2}$ and $Q=Q_{1}\oplus Q_{2}$. Then
$P\oplus Q^{(\omega)}\cong P_{1}\oplus P_{2}\oplus Q_{1}^{(\omega)}\oplus Q_{2}^{(\omega)}\cong P_{1}\oplus Q_{1}^{(\omega)}\oplus P_{2}\oplus Q_{2}^{(\omega)}\cong P_{1}\oplus P_{2}=P.$
On the other hand, by Lemma 1.5, $J=\operatorname{Tr}_{\Lambda}(Q)=\operatorname{Tr}_{\Lambda}(Q_{1})+\operatorname{Tr}_{\Lambda}(Q_{2})=J_{1}+J_{2}$, so
$P/PJ\cong\frac{P_{1}}{P_{1}(J_{1}+J_{2})}\oplus\frac{P_{2}}{P_{2}(J_{1}+J_{2})}\cong\frac{P_{1}/P_{1}J_{1}}{P_{1}(J_{1}+J_{2})/P_{1}J_{1}}\oplus\frac{P_{2}/P_{2}J_{2}}{P_{2}(J_{1}+J_{2})/P_{2}J_{2}}$
and $P/PJ$ is a finitely generated projective right $\Lambda/J$-module because it is a direct sum of quotients of finitely generated projective modules. Therefore $[P]\in B(\Lambda)$.
$(ii)$ is proved similarly.
Lemma 2.5.
Let $\Lambda$ be a ring, and let $I\in\mathcal{T}(\Lambda)$. Then there is a semigroup embedding $V(\Lambda/I)\hookrightarrow V(\Lambda)\sqcup B(\Lambda)$. [^1] If $I\neq 0$, it cannot be made into a monoid embedding because $[0]$ maps to the isomorphism class of a non-finitely generated projective module.
Proof.  Let $Q$ be a countably generated projective right $\Lambda$-module with trace ideal $I$. Let $P^{\prime}$ be a projective right $\Lambda/I$-module. By Lemma 1.9, there exists a projective right $\Lambda$-module $P$ such that $I\subseteq\operatorname{Tr}_{\Lambda}(P)$, $P/PI\cong P^{\prime}$ and $\operatorname{Tr}_{\Lambda}(P)/I=\operatorname{Tr}_{\Lambda/I}(P^{\prime})$.
Note that $P\oplus Q^{(\omega)}$ is $I$-big, since $P\oplus Q^{(\omega)}\cong P\oplus Q^{(\omega)}\oplus Q^{(\omega)}$. By Lemma 1.5, $QI=Q$, so $(P\oplus Q^{(\omega)})/(P\oplus Q^{(\omega)})I\cong P/PI$ is finitely generated. Therefore, $[P\oplus Q^{(\omega)}]\in V(\Lambda)\sqcup B(\Lambda)$. The isomorphism class of $P\oplus Q^{(\omega)}$ is independent on the choice of $Q$: let $Q^{\prime}$ be another countably generated projective right $\Lambda$-module with trace ideal $I$. Then, by Lemma 1.5, $Q^{\prime}I=Q^{\prime}$, so $(P\oplus Q^{\prime(\omega)})/(P\oplus Q^{\prime(\omega)})I\cong P/PI$. Hence, by Lemma 1.14, $[P\oplus Q^{(\omega)}]=[P\oplus Q^{\prime(\omega)}]$.
Let $\iota_{I}\colon V(\Lambda/I)\longrightarrow V(\Lambda)\sqcup B(\Lambda)$ be the map defined by $\iota_{I}[P^{\prime}]=[P\oplus Q^{(\omega)}]$. It is easy to check that this is well-defined: if $[P_{1}^{\prime}]=[P_{2}^{\prime}]$, then $P_{1}/P_{1}I\cong P_{2}/P_{2}I$. Since $P_{i}\oplus Q^{(\omega)}$ is $I$-big and $(P_{i}\oplus Q^{(\omega)})/(P_{i}\oplus Q^{(\omega)})I\cong P_{i}/P_{i}I$ for each $i=1,2$, by Lemma 1.14, $[P_{1}\oplus Q^{(\omega)}]=[P_{2}\oplus Q^{(\omega)}]$ and then $\iota_{I}[P_{1}^{\prime}]=\iota_{I}[P_{2}^{\prime}]$. It is easy to see that it is a semigroup homomorphism:
$\iota_{I}[P_{1}^{\prime}\oplus P_{2}^{\prime}]$ $=[P_{1}\oplus P_{2}\oplus Q^{(\omega)}]=[P_{1}\oplus Q^{(\omega)}\oplus P_{2}\oplus Q^{(\omega)}]$
$=[P_{1}\oplus Q^{(\omega)}]+[P_{2}\oplus Q^{(\omega)}]=\iota_{I}[P_{1}^{\prime}]+\iota_{I}[P_{2}^{\prime}].$
Finally, to see that it is injective, let $[P_{1}^{\prime}],[P_{2}^{\prime}]\in V(\Lambda/I)$ such that $\iota_{I}[P_{1}^{\prime}]=\iota_{I}[P_{2}^{\prime}]$. Then $[P_{1}\oplus Q^{(\omega)}]=[P_{2}\oplus Q^{(\omega)}]$, so $P_{1}\oplus Q^{(\omega)}\cong P_{2}\oplus Q^{(\omega)}$. By Lemma 1.5, $QI=Q$, so $(P_{i}\oplus Q^{(\omega)})/(P_{i}\oplus Q^{(\omega)})I\cong P_{i}/P_{i}I\cong P_{i}^{\prime}$ for each $i=1,2$, and hence $[P_{1}^{\prime}]=[P_{2}^{\prime}]$.
Remark 2.6.
Consider the set $\mathcal{T}(\Lambda)$ and the disjoint union $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$. We will denote the elements of $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$ by $([P],I)$ with $[P]\in V(\Lambda/I)$. Note that, if $I\subseteq J$, there is a morphism of monoids $f_{IJ}\colon V(\Lambda/I)\to V(\Lambda/J)$ defined by $f_{IJ}([X],I)=([X/XJ],J)$ with $[X]\in V(\Lambda/I)$. Therefore, $f_{IJ}(V(\Lambda/I))\subseteq V(\Lambda/J)$. Define addition in $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$ by
$([P_{1}],I)+([P_{2}],J)$ $=f_{I,I+J}([P_{1}],I)+f_{J,I+J}([P_{2}],J)$
$=([P_{1}/P_{1}J],I+J)+([P_{2}/P_{2}I],I+J)$
$=([P_{1}/P_{1}J\oplus P_{2}/P_{2}I],I+J)$
for each $[P_{1}]\in V(\Lambda/I)$, $[P_{2}]\in V(\Lambda/J)$. Therefore, the universal property of the disjoint union and Lemma 2.5 implies that there exists a semigroup embedding
$\iota\colon\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)\to V(\Lambda)\sqcup B(\Lambda).$
The following lemma states that this embedding is precisely an isomorphism.
Corollary 2.7.
[wiegand, Theorem 3.8] Let $\Lambda$ be a ring. Then
  1. (i)
    $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)\cong V(\Lambda)\sqcup B(\Lambda)$ as monoids.
  2. (ii)
    $\bigsqcup_{I\in\mathcal{T}(\Lambda)\setminus\{0\}}V(\Lambda/I)\cong B(\Lambda)$ as semigroups.
Proof.  The isomorphism in $(i)$ is given by $\varphi([X],I)=\iota_{I}([X])$, where $\iota_{I}$ is the embedding from Lemma 2.5. To prove $(ii)$, note that elements in $V(\Lambda)$ can be thought of as elements in $V(\Lambda/I)$ for $I=\{0\}$, so $(ii)$ follows from $(i)$.
The following result gives a characterization of when every element in $B(\Lambda)$ is in the same isomorphism class as a direct sum of finitely generated projective modules.
Proposition 2.8.
Let $\Lambda$ be a ring. Then $B(\Lambda)$ is a semigroup of modules such that any representative is isomorphic to a direct sum of finitely generated projective modules if and only if the following two properties are satisfied:
  1. (1)
    If $I$ is the trace ideal of a projective right $\Lambda$-module, then there is a right $\Lambda$-module $P$, which is a direct sum of finitely generated projective modules such that $\operatorname{Tr}_{\Lambda}(P)=I$.
  2. (2)
    If $I$ is the trace ideal of a countably generated projective right $\Lambda$-module and $P^{\prime}$ is a finitely generated projective right $\Lambda/I$-module, then there is a finitely generated projective right $\Lambda$-module $X$ such that $P^{\prime}\cong X/XI$.
Proof.  Assume that (1) and (2) are satisfied, and let $[P]\in B(\Lambda)$. Then there exists a countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ such that $P$ is $J$-big and $P/PJ$ is a finitely generated projective right $\Lambda/J$-module.
By (1), $Q$ can be taken to be a direct sum of finitely generated projective modules. By (2) applied to $P/PJ$, there exists a finitely generated projective right $\Lambda$-module $X$ such that $X/XJ\cong P/PJ$. Let $P^{\prime}=X\oplus Q^{(\omega)}$. Clearly, $[P^{\prime}]\in B(\Lambda)$, and by Lemma 1.14, $[P]=[P^{\prime}]$.
Assume now that every element in $B(\Lambda)$ is in the same isomorphic class as a direct sum of finitely generated projective modules. Let $P$ be a projective right $\Lambda$-module with trace $I$. By [kaplansky], $P=\bigoplus_{i\in J}P_{i}$ is a direct sum of countably generated projective modules. By Example 2.3, $[P_{i}^{(\omega)}]\in B(\Lambda)$. Therefore, $P_{i}^{(\omega)}$ is isomorphic to a direct sum of finitely generated projective modules for each $i\in J$. Hence, $P^{(\omega)}$ is isomorphic to a direct sum of finitely generated projective modules and $I=\operatorname{Tr}_{\Lambda}(P^{(\omega)})$. This gives (1).
To prove (2), let $Q$ be a countably generated projective right $\Lambda$-module with trace $I$, and let $P^{\prime}$ be a finitely generated projective right $\Lambda/I$-module. By Lemma 1.9, since $I$ is the trace of a countably generated projective module, there exists a countably generated projective right $\Lambda$-module $P$ such that $I\subseteq\operatorname{Tr}_{\Lambda}(P)$, $P/PI\cong P^{\prime}$ and $\operatorname{Tr}_{\Lambda}(P)/I=\operatorname{Tr}_{\Lambda/I}(P^{\prime})$.
Now, $[P\oplus Q^{(\omega)}]\in B(\Lambda)$, so it is a direct sum of finitely generated projective modules. Write $P\oplus Q^{(\omega)}=\bigoplus_{i\in K}X_{i}$ with $X_{i}$ a finitely generated projective $\Lambda$-module for each $i\in K$. Since $(\bigoplus_{i\in K}X_{i})/(\bigoplus_{i\in K}X_{i})I\cong P^{\prime}$ is finitely generated, there is a finite subset $K_{0}\subseteq K$ such that $(\bigoplus_{i\in K_{0}}X_{i})/(\bigoplus_{i\in K_{0}}X_{i})I\cong P^{\prime}$, so $X=\bigoplus_{i\in K_{0}}X_{i}$ is the finitely generated projective $\Lambda$-module we were looking for.
Remark 2.9.
Let $\Lambda$ be a ring. It is well known that the functor $(-)^{*}=\mathrm{Hom}_{\Lambda}(-,\Lambda)$ induces a duality between finitely generated projective right $\Lambda$-modules and finitely generated projective left $\Lambda$-modules. So that, $\varphi\colon V_{r}(\Lambda)\to V_{\ell}(\Lambda)$ defined by $\varphi[P]=[P^{*}]$, is an isomorphism of monoids. Moreover, $P$ and $P^{*}$ have the same trace (cf. [andersonfuller, Exercise 22.1]).
In the next example, we show that the canonical monoid isomorphism $V_{r}(\Lambda)\cong V_{\ell}(\Lambda)$ of Remark 2.9, in general, does not extend to an isomorphism between the monoids $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$. We will show that it does extend, provided that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$.
Example 2.10.
In [TAMS, Example 3.6 (iv)] it is given an example of a semilocal ring $\Lambda$ such that $\Lambda/J(\Lambda)\cong D_{1}\times D_{2}$, with $D_{1}$ and $D_{2}$ division rings, such that any projective left $\Lambda$-module is free but this is not true on the right. Though, on the right, it is still true that $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$.
In such $\Lambda$, what happens is that one of the maximal ideals contains a nonzero trace of a countably generated projective right $R$-module. Therefore $\mathcal{T}_{\ell}(\Lambda)=\{0\},\Lambda\subsetneq\mathcal{T}_{r}(\Lambda)$. Then $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$ are not isomorphic. This cannot happen if $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$.
Assume $\Lambda$ is a ring such that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$. In view of the operation defined in Remark 2.6, to define a monoid homomorphism from $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ to $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$, that extends the isomorphism between the monoids of finitely generated projective modules of Remark 2.9, we have to be careful with the compatibility of the operations defined in both monoids. That is, identifying $[P]\in V_{r}(\Lambda/I)$ with $([P],I)\in V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and defining the map by $([P],I)\mapsto([P^{*}],I)$, then we must check that $([P_{1}],I)+([P_{2}],J)$ is mapped to $([P_{1}^{*}],I)+([P_{2}^{*}],J)$.
The following result is a technical lemma that is crucial for our discussion.
Lemma 2.11.
Let $\Lambda$ be a ring, and let $I,J$ be ideals of $\Lambda$. Let $P$ be a finitely generated projective right $\Lambda/I$-module. Then
$\operatorname{Hom}_{\Lambda/(I+J)}(P/PJ,\Lambda/(I+J))\cong\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)/((I+J)/I\cdot\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)).$
Proof.  Note that we can identify $P/PJ$ with $P\otimes_{\Lambda/I}\Lambda/(I+J)$. Then
$\operatorname{Hom}_{\Lambda/(I+J)}(P/PJ,\Lambda/(I+J))$ $\cong\operatorname{Hom}_{\Lambda/(I+J)}(P\otimes_{\Lambda/I}\Lambda/(I+J),\Lambda/(I+J))$
$\cong\operatorname{Hom}_{\Lambda/I}(P,\operatorname{Hom}_{\Lambda/(I+J)}(\Lambda/(I+J),\Lambda/(I+J)))$
$\cong\operatorname{Hom}_{\Lambda/I}(P,\Lambda/(I+J)).$
where in the second isomorphism we have used the Hom-tensor adjunction. Now, we want to prove that $\operatorname{Hom}_{\Lambda/I}(P,\Lambda/(I+J))\cong\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)/J\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$ as left $\Lambda/I$-modules.
Consider the map $\varphi\colon\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\to\operatorname{Hom}_{\Lambda/I}(P,\Lambda/(I+J))$ defined by $\varphi(f)=\pi\circ f$, where $\pi\colon\Lambda/I\to\Lambda/(I+J)$ is the projection modulo $(I+J)/I$. Clearly, it is a homomorphism of left $\Lambda/I$-modules. Let $g\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/(I+J))$. Since $P$ is projective, there is a lift $G\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$ making the following diagram commute
${P}$${\Lambda/I}$${\Lambda/(I+J)}$00{0}$\scriptstyle{g}$$\scriptstyle{G}$$\scriptstyle{\pi}$
Therefore, $\varphi$ is surjective. The kernel of this homomorphism is
$\ker\varphi$ =$=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid\pi\circ f=0\}$
=$=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid f(x)\in\ker\pi\text{ for every }x\in P\}$
=$=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid f(x)\in(I+J)/I\text{ for every }x\in P\}$
$=\operatorname{Hom}_{\Lambda/I}(P,(I+J)/I).$
It remains to prove that $\operatorname{Hom}_{\Lambda/I}(P,(I+J)/I)=(I+J)/I\cdot\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$.
Let $f\in\operatorname{Hom}_{\Lambda/I}(P,(I+J)/I)$. Consider $(I+J)/I$ as a right $\Lambda/I$-module. Then there exists an epimorphism $\pi\colon(\Lambda/I)^{(A)}\to(I+J)/I$ for some set $A$. Since $P$ is projective, there exists a lift $F\in\operatorname{Hom}_{\Lambda/I}(P,(\Lambda/I)^{(A)})$ such that $\pi\circ F=f$, that is, the following diagram commutes
${P}$${(\Lambda/I)^{(A)}}$${(I+J)/I}$${0.}$$\scriptstyle{f}$$\scriptstyle{F}$$\scriptstyle{\pi}$
Since $P$ is finitely generated, the image of $F$ is contained in a finite number of copies of $\Lambda/I$, say $(\Lambda/I)^{(A_{0})}$, with $A_{0}\subseteq A$ finite. Let $\{x_{i}\}_{i\in A}$ be a basis of $(\Lambda/I)^{(A)}$. Then
$f(x)=\pi\circ F(x)=\pi\Big{(}\sum_{i\in A_{0}}x_{i}a_{i}\Big{)}=\sum_{i\in A_{0}}\pi(x_{i})a_{i}=\sum_{i\in A_{0}}\pi(x_{i})\cdot\varphi_{i}(F(x)).$
where $\varphi_{i}\colon(\Lambda/I)^{(A_{0})}\to\Lambda/I$ denotes the projection into the $i$th component, $i\in A_{0}$. Hence, $f=\sum_{i\in A_{0}}\pi(x_{i})\cdot(\varphi_{i}\circ F)\in(I+J)/I\cdot\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$.
Conversely, let $F=\sum_{i=1}^{n}(y_{i}+I)\cdot F_{i}\in(I+J)/I\cdot\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$ for some $y_{1},\dotsc,y_{n}\in J$ and $F_{1},\dotsc,F_{n}\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$. Then regarding the product of elements as composition of homomorphisms, $F\in\operatorname{Hom}_{\Lambda/I}(P,(I+J)/I)$. Therefore, by the First Isomorphism Theorem,
$\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)/((I+J)/I\cdot\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I))\cong\operatorname{Hom}_{\Lambda/I}(P,\Lambda/(I+J)),$
so the statement is proved.
Corollary 2.12.
Let $\Lambda$ be a ring such that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$. Then there is an isomorphism of monoids between $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$.
Proof.  Let $\varphi\colon V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\to V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$ be the map defined by $\varphi([P],I)=([P^{*}],I)$. By Remark 2.9, it is well-defined. By Lemma 2.11, it is a monoid homomorphism since
$\varphi(([P_{1}],I)+([P_{2}],J))$ $=\varphi([P_{1}/P_{1}J\oplus P_{2}/P_{2}I],I+J)$
$=([\operatorname{Hom}_{\Lambda/(I+J)}(P_{1}/P_{1}J\oplus P_{2}/P_{2}I,\Lambda/(I+J))],I+J)$
$=([\operatorname{Hom}_{\Lambda/(I+J)}(P_{1}/P_{1}J,\Lambda/(I+J))$
$\qquad\qquad\oplus\operatorname{Hom}_{\Lambda/(I+J)}(P_{2}/P_{2}I,\Lambda/(I+J))],I+J)$
$=([\operatorname{Hom}_{\Lambda/I}(P_{1},\Lambda/I)/J\operatorname{Hom}_{\Lambda/I}(P_{1},\Lambda/I))$
$\qquad\qquad\oplus\operatorname{Hom}_{\Lambda/J}(P_{2},\Lambda/J)/I\operatorname{Hom}_{\Lambda/J}(P_{2},\Lambda/J))],I+J)$
$=([\operatorname{Hom}_{\Lambda/I}(P_{1},\Lambda/I)],I)+([\operatorname{Hom}_{\Lambda/J}(P_{2},\Lambda/J)],J)$
$=\varphi([P_{1},I])+\varphi([P_{2}],J)$
Finally, by Remark 2.9 it is bijective.
Then Theorem 1.18 together with the fact that, over a noetherian ring, $\Lambda$, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ is just the set of idempotent ideals of $\Lambda$ (cf. [whitehead]), immediately yields.
Theorem 2.13.
[P2, Theorem 2.12] Let $\Lambda$ be a noetherian ring satisfying condition ()(*)( ∗ ). Then $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\cong V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$.
In general, it is an open question whether $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ for $\Lambda$ a right noetherian ring (cf. Corollary 4.9 and the comments after it for further information around this question).
3. Lifting projectives modules modulo a right $T$-nilpotent ideal
Lemma 3.1.
Let $\Lambda$ be a ring, and let $N$ be a right $T$-nilpotent ideal of $\Lambda$.
  1. (1)
    Let $X$ be a right projective $\Lambda/N$-module. Then there exists a projective right $\Lambda$-module $P$ such that $P/PN\cong X$. All projective modules with this property are isomorphic. In addition, $X$ is finitely generated or countably generated if and only if so is $P$.
  2. (2)
    Let $I$ be an idempotent ideal of $\Lambda$ such that $(I+N)/N$ is the trace of a countably generated projective right $\Lambda/N$-module $X$. Then $I$ is the trace of any countably generated projective right $\Lambda$-module $P$ such that $P/PN\cong X$.
Proof.  $(1)$ Let $X$ be a projective right $\Lambda/N$-module. As $X$ is a direct sum of countably generated projective right $\Lambda/N$-modules, we may assume that $X$ is countably generated. In this case, $X$ is the direct limit of a direct system
[…]
of finitely generated free $\Lambda/N$-modules, and the canonical presentation of the direct limit
[…]
splits. We can lift (1) to a countable direct system of finitely generated free $\Lambda$-modules
[…]
such that when we apply the functor $-\otimes_{\Lambda}\Lambda/N$ to (3) we obtain a direct system isomorphic to (1).
Let $P$ be the direct limit of (3), and consider its canonical presentation
[…]
As (2) splits, there exists $\Psi\colon\bigoplus_{n\geq 1}{F}_{n}\to\bigoplus_{n\geq 1}{F}_{n}$ such that $\Psi\circ\Phi=\mathrm{Id}+B$ where $B$ can be thought of as a column finite matrix with entries in $N$ (that is, $B\colon\bigoplus_{n\geq 1}{F}_{n}\to\bigoplus_{n\geq 1}{F}_{n}$ has the image contained in $\bigoplus_{n\geq 1}{F}_{n}N$). Since $N$ is right $T$-nilpotent, $B$ is in the Jacobson radical of the ring of countable column-finite matrices with coefficients in $\Lambda$ [SW], and then $\mathrm{Id}+B$ is invertible. So that $(\mathrm{Id}+B)^{-1}\circ\Psi$ is a left inverse of $\Phi$ and we can deduce that $P$ is projective.
This shows that any lifting of the direct system (1) gives a projective $\Lambda$-module $P$ that, in addition, satisfies $P/PN\cong X$. As $PN$ is a small submodule of $P$ by [andersonfuller, Lemma 28.3], $P/PN\cong X$ is finitely generated (countably generated) if and only if so is $P$.
$(2)$. By $(1)$, the module $X$ can be lifted to a countably generated projective right $\Lambda$-module $P$ such that $P/PN\cong X$. Let $J$ be the trace of $P$, then $J+N=I+N$. By [traces, Corollary 2.9], $J\subseteq I$ because $J$ is a trace ideal of a projective module and $N$, being right $T$-nilpotent, is contained in the Jacobson radical of $\Lambda$. Since $J+N=I+N$ and $I^{2}=I$, $IJ+IN=I$. By [andersonfuller, Lemma 28.3], $IN\ll I$ which implies $IJ=I$ and, in particular, $I\subseteq J$. Therefore, $I=J$ is the trace of $P$.
Lemma 3.2.
Let $\Lambda$ be a ring, and let $N$ be a right $T$-nilpotent ideal of $\Lambda$. Let $I$ be an ideal of $\Lambda$, and let $P$ be a countably generated right $\Lambda$-module. Then, $P$ is $I$-big if and only if $P/PN$ is $(I+N)/N$-big.
Proof.  Assume that $P$ is $I$-big. Let $X$ be a countably generated projective right $\Lambda/N$-module such that its trace is contained in $(I+N)/N$. By Lemma 3.1, there exists a countably generated projective right $\Lambda$-module $Q$ such that $Q/QN\cong X$; therefore its trace $J$ is contained in $I+N$. Since $J$ is the trace of a countably generated projective, and $N$ is right $T$-nilpotent, we deduce from [traces, Corollary 2.9] that $J\subseteq I$. By hypothesis, there exists an onto module homomorphism $f\colon P\to Q$. Therefore, there is an onto module homomorphism
${g\colon P}$${Q}$${Q/QN\cong X}$$\scriptstyle{f}$$\scriptstyle{\pi}$
where $\pi$ denotes the canonical projection. As $PN$ is contained in the kernel of $g$, the morphism $g$ factors through an onto homomorphism $P/PN\to X$. This shows that $P/PN$ is $(I+N)/N$-big.
To prove the converse, let $Q$ be a countably generated projective right $\Lambda$-module such that its trace is contained in $I$. Then the trace of $Q/QN$ is contained in $I+N/N$ and since $P/PN$ is $(I+N)/N$-big, there exists an onto module homomorphism $g\colon P/PN\to Q/QN$. Since $P$ is projective, there exists $f\colon P\to Q$ such that the diagram
${P}$${Q}$${Q/QN}$$\scriptstyle{f}$$\scriptstyle{g\circ\pi_{1}}$$\scriptstyle{\pi_{2}}$
is commutative, and where $\pi_{1}\colon P\to P/PN$ and $\pi_{2}\colon Q\to Q/QN$ denote canonical projections. Therefore, $Q=f(P)+QN$. By [andersonfuller, Lemma 28.3], $QN$ is a small submodule of $Q$, and we deduce that $f(P)=Q$. This shows that $P$ is $I$-big.
Corollary 3.3.
Let $\Lambda$ be a ring, and let $N$ be a right $T$-nilpotent ideal of $\Lambda$. Then the functor $-\otimes_{\Lambda}\Lambda/N$ induces a monoid isomorphism $\varphi\colon V^{*}(\Lambda)\to V^{*}(\Lambda/N)$ which restricts to monoid isomorphism $V(\Lambda)\to V(\Lambda/N)$ and $V(\Lambda)\sqcup B(\Lambda)\to V(\Lambda/N)\sqcup B(\Lambda/N)$.
Proof.  Consider the monoid morphism $\varphi\colon V^{*}(\Lambda)\to V^{*}(\Lambda/N)$ defined by $\varphi([P])=[P/PN]$ where $P$ denotes a countably generated, projective, right $\Lambda$-module. It follows from Lemma 3.1 that $\varphi$ is an isomorphism that restricts to an isomorphism $V(\Lambda)\to V(\Lambda/N)$.
Let $P$ be a countably generated projective $\Lambda$-module such that $[P]\in B(\Lambda)$. That is there exists $I$, the trace of a nonzero countably generated projective module $Q$, such that $P$ is $I$-big and $P/PI$ is finitely generated.
By Lemma 1.6, $I+N/N$ is the trace of $Q/QN$. Notice also that $I+N/N\neq\{0\}$. By Lemma 3.2, $P/PN$ is $I+N/N$-big. Clearly
$(P/PN)/(P(I+N)/PN)\cong P/P(I+N)$
is finitely generated. Hence $\varphi([P])\in B(\Lambda/N)$.
Let $X$, $Y$ be countably generated projective right $\Lambda/N$-modules. Let $K$ be the trace of $Y$ in $\Lambda/N$. Assume that $X$ is $K$-big and that $X/XK$ is finitely generated.
By Lemma 3.1 there exist countably generated projective $\Lambda$-modules $P$ and $Q$ such that $P/PN\cong X$ and $Q/QN\cong Y$. Let $I$ be the trace of $Q$. By Lemma 1.6, $K=I+N/N$.
By Lemma 3.2, $P$ is $I$-big. Moreover, $P/PI$ is finitely generated because so is $P/P(I+N)$. This shows that $\varphi^{-1}([X])\in B(\Lambda)$. This concludes the proof that $\varphi$ induces an isomorphism $V(\Lambda)\sqcup B(\Lambda)\to V(\Lambda/N)\sqcup B(\Lambda/N)$.
4. The case of right noetherian PI-rings
In this section we shall prove that countably generated, projective, right or left modules over a right noetherian PI-ring are always relatively big. The key result will be to show that they satisfy the ()(*)( ∗ ) condition introduced in §2. As a right noetherian PI-ring $R$ has nilpotent nilradical $N(R)$ and $R/N(R)$ is two-sided noetherian, the result will follow by combining Theorem 1.18 with the results in §3.
First we consider the case of commutative noetherian rings. P. Příhoda and G. Puninski proved in [PP] that commutative noetherian rings satisfy condition ()(*)( ∗ ). Here we give another proof of this result, which we find interesting as it is quite different from the one given in [PP].
4.1. Commutative noetherian rings
[…]
Lemma 4.1.
Let $\Lambda$ be a ring. Let $(a_{k})_{k\geq 1}$ be a sequence of elements of $\Lambda$ satisfying $a_{k}a_{k+1}=a_{k+1}$ for any $k\geq 1$. Set $b_{k}=1-a_{k}$ for any $k\geq 1$. Then,
  1. (i)
    there is an descending sequence of principal right ideals
    $a_{1}\Lambda\supseteq a_{2}\Lambda\supseteq\cdots\supseteq a_{k}\Lambda\supseteq\cdots$ (4)
  2. (ii)
    $b_{k}b_{k+1}=b_{k}$ for any $k\geq 1$, so there is an ascending sequence of principal left ideals
    $\Lambda b_{1}\subseteq\Lambda b_{2}\subseteq\cdots\subseteq\Lambda b_{k}\subseteq\cdots$ (5)
  3. (iii)
    For any $k\geq 1$,
    $\Lambda b_{k}\subseteq\mathrm{l.ann}(a_{k+1})\subseteq\Lambda b_{k+1}$
    and
    $a_{k}\Lambda\supseteq\mathrm{r.ann}(b_{k})\supseteq a_{k+1}\Lambda.$
  4. (iv)
    $I=\bigcup_{k\geq 1}\Lambda b_{k}$ is a left pure ideal of $R$ and it is also projective.
  5. (v)
    the sequence (4) is stationary if and only if so is the sequence (5). In this case, there exists $k_{0}$ such that $b_{k_{0}}$ and $a_{k_{0}}$ are idempotent, $I=\Lambda b_{k_{0}}$ and $\bigcap_{k\geq 1}a_{k}\Lambda=a_{k_{0}}\Lambda$.
Proof.  Statements $(i)$, $(ii)$ and $(iii)$ follow easily by direct computation.
The purity claimed in statement $(iv)$ was proved by Lazard in [lazard, Lemme 2]. Then, as $\Lambda/I$ is a countably presented flat left $\Lambda$-module, $I$ is projective. For a self-contained proof, the reader can check [FHS1, Lemma 2].
Statement $(iii)$ readily implies that (4) is stationary if and only if so is (5).
If $a_{k}\Lambda=a_{k+1}\Lambda$, then there exists $\lambda\in\Lambda$ such that $a_{k+1}\lambda=a_{k}$. Hence, $a_{k+1}\lambda a_{k+1}=a_{k+1}$ which implies that $a_{k}=a_{k+1}\lambda$ is idempotent.
If $\Lambda b_{k}=\Lambda b_{k+1}$, then there exists $\mu\in\Lambda$ such that $\mu b_{k}=b_{k+1}$. As before, this implies that $b_{k+1}$ is idempotent.
Assume the sequences (4) and (5) are stationary. Then the argument above shows that, taking $k_{0}$ large enough, the second part of statement $(v)$ holds.
Lemma 4.2.
Let $R$ be a commutative ring. Let $I\supseteq J$ be ideals of $R$ such that $JI=J$. If $J$ is finitely generated, then there exists $a\in I$ such that $xa=x$ for any $x\in J$. In particular $I\supseteq aR\supseteq J$.
Proof.  Let $J=x_{1}R+\dots+x_{n}R$. The equality $JI=J$ implies that there exists $A\in M_{n}(I)$ such that $(x_{1},\dots,x_{n})A=(x_{1},\dots,x_{n})$ or, equivalently
$(x_{1},\dots,x_{n})(\mathrm{Id}-A)=(0,\dots,0).$
Multiplying the above identity on the right by the adjoint matrix of $\mathrm{Id}-A$, we deduce that $\det(A)$ annihilates $J$. To conclude the proof, note that there exists $a\in I$ such that $\det(A)=1-a$.
Proposition 4.3.
([PP, Lemma 2.1]) Every commutative noetherian ring satisfies condition ()(*)( ∗ ). This is to say that if $R$ is a commutative noetherian ring and
$J_{1}\supseteq J_{2}\supseteq\cdots\supseteq J_{k}\supseteq\cdots$ (6)
is a descending chain of ideals of $R$ satisfying $J_{k+1}J_{k}=J_{k+1}$ for each $k\geq 1$, then (6) is stationary.
In addition, in this context, $\bigcap_{k\geq 1}J_{k}=eR$ where $e$ is an idempotent of $R$.
Proof.  Let
$J_{1}\supseteq J_{2}\supseteq\cdots\supseteq J_{k}\supseteq\cdots$ (*)
be a descending chain of ideals of $R$ such that $J_{k+1}J_{k}=J_{k+1}$, for each $k\geq 1$.
Since $R$ is noetherian, the ideals $J_{k}$ are finitely generated for every $k\geq 1$. By Lemma 4.2, there exists a sequence of elements $(a_{k})_{k\geq 1}$ in $R$ such that
$J_{1}\supseteq a_{1}R\supseteq J_{2}\supseteq a_{2}R\supseteq\cdots\supseteq J_{k}\supseteq a_{k}R\supseteq\cdots$
and such that $a_{k+1}a_{k}=a_{k+1}$.
By Lemma 4.1 and because $R$ is noetherian, the ascending chain of ideals
$(1-a_{1})R\subseteq(1-a_{2})R\subseteq\cdots\subseteq(1-a_{k})R\subseteq\cdots$
is stationary and its union is generated by an idempotent of $R$. By Lemma 4.1, so is the descending sequence
$a_{1}R\supseteq a_{2}R\supseteq\cdots\supseteq a_{k}R\supseteq\cdots$
and its intersection is generated by an idempotent of $R$. The sequence (*4.1) is also stationary and the intersection of these ideals is generated by an idempotent of $R$.
In the following result, we recover a result of Bass that shows that, over commutative noetherian rings, the theory of countably generated projective modules reduces to the finitely generated ones.
Corollary 4.4.
[bass] Let $R$ be a commutative noetherian ring. Let $P$ be a countably generated projective module that is not finitely generated. Then there exists an idempotent $e\in R$ with $P\cong(1-e)P\oplus\left(eR\right)^{(\omega)}$ and such that $(1-e)P$ is finitely generated.
Proof.  Let $P$ be a countably generated projective module over $R$. By Proposition 4.3 and Theorem 1.18, there exists an idempotent $e\in R$ such that $P/eP$ is finitely generated, $I=eR$ is an ideal minimal with respect to that property, and $P$ is $I$-big. Therefore $P\cong P\oplus(eR)^{(\omega)}$.
Since $R$ is commutative
$P=eP\oplus(1-e)P\cong eP\oplus(1-e)P\oplus(eR)^{(\omega)}\cong(1-e)P\oplus(eR)^{(\omega)},$
where the last isomorphism follows by the Eilenberg swindle. To conclude the proof of the statement, observe that $(1-e)P\cong P/eP$.
Corollary 4.4 implies that a countably generated projective module over a commutative noetherian domain that is not finitely generated is free.
Not much is known on the structure of infinitely generated projective modules over general commutative domains. To our knowledge, the basic questions on the topic posed in [fuchssalce, p.  246] are still widely open and it seems that to answer them one needs to go beyond the theory of relatively big projectives.
In general, over a commutative domain $R$, the only trace ideals are $\{0\}$ and $R$, so $B(R)$ contains only the isomorphism class of $R^{(\omega)}$. Therefore, $V(R)\sqcup B(R)=V(R)\sqcup\{[R^{(\omega)}]\}$. We do not know any reasonable characterization of domains $R$ such that $V^{*}(R)=V(R)\sqcup B(R)$, neither examples where this fails.
4.2. One-sided noetherian PI-rings
By the results in [P3] or in [wiegand] that the structure of countably projective modules over finitely generated algebras over a commutative noetherian ring is more involved than the one described in Corollary 4.4. The results on these papers were depending on the fact that if $\Lambda$ is a finitely generated algebra over a commutative noetherian ring of Krull dimension $1$ then $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$, that is, all countably generated projective, right or left, $\Lambda$-modules are relatively big. In this subsection we show this is true for countably generated projective modules over any right noetherian, PI-ring $\Lambda$.
Lemma 4.5.
Let $\Lambda$ be a ring. Let $I\supseteq J$ be ideals of $\Lambda$ such that either $JI=J$ or $IJ=J$. Let $K_{1},\dots,K_{n}$ be ideals of $\Lambda$ (not necessarily different) such that $K_{1}\cdots K_{n}\subseteq J$ and $J+K_{i}=I+K_{i}$, for $i=1,\dots,n$. Then $J=J^{2}$.
Proof.  Using the hypothesis we deduce that,
$I^{n}\subseteq\prod_{i=1}^{n}(I+K_{i})=\prod_{i=1}^{n}(J+K_{i})\subseteq J.$
Assume $JI=J$,then $JI^{n}=J$. Since $I^{n}\subseteq J$, we deduce that $J\subseteq J^{2}$. This shows that $J=J^{2}$.
A symmetric argument gives the conclusion if $IJ=J$.
Proposition 4.6.
Let $\Lambda$ be a right noetherian ring. Let
$I_{1}\supseteq I_{2}\supseteq\cdots\supseteq I_{k}\supseteq\cdots$ (7)
be a sequence of ideals of $\Lambda$ satisfying that $I_{k+1}I_{k}=I_{k+1}$ for any $k\geq 1$. Let
$J_{1}\supseteq J_{2}\supseteq\cdots\supseteq J_{k}\supseteq\cdots$ (8)
be a sequence of ideals of $\Lambda$ satisfying that $J_{k}J_{k+1}=J_{k+1}$ for any $k\geq 1$. Then,
  1. (i)
    Let $N(\Lambda)$ denote the prime radical of $\Lambda$ (that is, the intersection of all prime ideals of $\Lambda$). If (7) or (8) are stationary in $\Lambda/N(\Lambda)$ then they are also stationary over $\Lambda$.
  2. (ii)
    Assume that $\Lambda$ is semiprime, and that (7) is stationary over $\Lambda/P$ for any minimal prime ideal $P$ of $\Lambda$. Then there exists $k_{0}\geq 1$ such that $I_{k_{0}+\ell}^{2}=I_{k_{0}+\ell}$ for any $\ell\geq 0$.
  3. (iii)
    Assume that $\Lambda$ is semiprime, and that (8) is stationary over $\Lambda/P$ for any minimal prime ideal $P$ of $\Lambda$. Then there exists $k_{0}\geq 1$ such that $J_{k_{0}+\ell}^{2}=J_{k_{0}+\ell}$ for any $\ell\geq 0$.
Proof.  $(i)$. By [GW, Theorem 2.11], $N(\Lambda)$ is nilpotent. Assume that (7) is stationary over $\Lambda/N(\Lambda)$. Then there exists $k_{1}\geq 0$ such that $I_{k_{1}}+N(\Lambda)=I_{k_{1}+\ell}+N(\Lambda)$ for any $\ell\geq 0$. By Lemma 4.5, this implies that $I_{k_{1}+1+\ell}^{2}=I_{k_{1}+1+\ell}$ for any $\ell\geq 0$. Since idempotent ideals that are equal modulo a nilpotent ideal must be equal, the claim follows.
A similar argument works if (8) is stationary.
$(ii)$. Since $\Lambda$ is a semiprime right Goldie ring, it has a finite number of minimal prime ideals $P_{1},\dots,P_{n}$, and its product is zero [GW, Proposition 6.1].
By hypothesis, there exists $k_{1}$ such that $I_{k_{1}+\ell}+P_{i}=I_{k_{1}}+P_{i}$ for any $\ell\geq 0$ and any $i=1,\dots,n$.
As in $(i)$, we use Lemma 4.5 to deduce that $k_{0}=k_{1}+1$ satisfies the conclusion of the statement.
Statement $(iii)$ is proved in a similar way.
Robson and Small proved in [RS] that PI-rings satisfying the ascending chain condition on two-sided ideals have a finite number of idempotent ideals.
We shall prove in Theorem 4.8 that, in addition, right noetherian PI-rings satisfy condition ()(*)( ∗ ) and its left analogue that we call ()(**)( ∗ ∗ ). The proof is patterned on the one by Robson and Small.
First we give some basic definitions needed to follow the proof.
Definition 4.7.
A ring $\Lambda$ is said to be a polynomial identity ring, a PI-ring for short, if there exists $n$ such that $\Lambda$ satisfies a monic polynomial of $\mathbf{Z}\langle x_{1},\dots,x_{n}\rangle$. That is, there exists a polynomial in non-commuting variables $x_{1},\dots,x_{n}$ with coefficients in $\mathbf{Z}$, $f(x_{1},\dots,x_{n})$ say, such that:
  • (1)
    $f(r_{1},\dots,r_{n})=0$ for any $r_{1},\dots,r_{n}\in\Lambda$,
  • (2)
    the coefficient of one of the monomials of highest degree of $f$ is one.
The polynomials in $\mathbf{Z}\langle x_{1},\dots,x_{n}\rangle$ of the form $\sum_{\sigma\in\Sigma_{n}}a(\sigma)x_{\sigma(1)}\cdots x_{\sigma(n)}$, where $\Sigma_{n}$ denotes the permutation group of $n$ elements and $a(\sigma)\in\mathbf{Z}$, are said to be multilinear polynomials or multilinear identities of degree $n$.
Of special importance in the theory is the standard identity of degree $n$
$s_{n}=s_{n}(x_{1},\dots,x_{n})=\sum_{\sigma\in\Sigma_{n}}\varepsilon(\sigma)x_{\sigma(1)}\cdots x_{\sigma(n)}$
where $\varepsilon(\sigma)$ denotes the sign of the permutation $\sigma$. For example, if $R$ is a non-zero commutative ring, a theorem of Amitsur and Levitzki shows that $M_{n}(R)$ satisfies the standard identity $s_{m}$ for any $m\geq 2n$ and does not satisfy $s_{2n-1}$, cf. [MR, §13.3.2].
A theorem of Kaplansky shows that if a ring satisfies a monic identity of degree at most $d$, then it satisfies a monic multilinear identity of degree at most $d$ (cf. [MR, §13.1.9]).
It is possible to define the PI-degree of a central simple algebra $\Lambda$ (that is, a simple algebra that is finite dimensional over its center) as the minimum $n$ such that $\Lambda$ satisfies $s_{2n}$ and not $s_{2n-1}$. As prime PI-rings localized at the non-zero elements of the center are central simple algebras, this definition of the PI-degree can be extended to prime PI-rings. For details, check [MR, §13.6.7].
Theorem 4.8.
Let $\Lambda$ be a right noetherian PI-ring. Then $\Lambda$ satisfies condition ()(*)( ∗ ) and its symmetric ()(**)( ∗ ∗ ).
Proof.  Assume the statement is not true. By noetherian induction, we may assume that $\Lambda$ is a right noetherian PI-ring that does not satisfy condition ()(*)( ∗ ) but so does any proper quotient ring of $\Lambda$.
By [RS, Theorem 4], $\Lambda$ has only a finite number of idempotent ideals. By Proposition 4.6 this implies that $\Lambda$ is a prime, right noetherian, PI-ring. Therefore, $\Lambda$ is (right and left) noetherian by a theorem of G. Cauchon [MR, 13.6.15 Theorem]. Now we adapt the steps of the proof of [RS, Theorem 4] to get a contradiction.
Let
$\Lambda\supsetneq I_{1}\supsetneq I_{2}\supsetneq\cdots\supsetneq I_{k}\supsetneq\cdots$ (9)
be a sequence of ideals of $\Lambda$ satisfying that $I_{k+1}I_{k}=I_{k+1}$ for any $k\geq 1$. Let $P$ be a prime ideal containing $I_{k}$ for some $k$ and assume that $\mathrm{deg}\,\Lambda/P=\mathrm{deg}\,\Lambda$. By [small, Theorem], the set of elements which are regular modulo $P$ consists of regular elements of $\Lambda$, forms an Ore set $S$ and the localization $\Lambda S^{-1}$ is a local ring with maximal ideal $P\Lambda S^{-1}$. Since $I_{k}$ is contained in $P$, so is $I_{k+\ell}$ for any $\ell\geq 1$. But in the local ring $\Lambda S^{-1}$, $P\Lambda S^{-1}I_{k+\ell}\Lambda S^{-1}=I_{k+\ell}\Lambda S^{-1}$ for any $\ell\geq 1$. As the ring is noetherian, Nakayama’s Lemma implies that $I_{k+\ell}\Lambda S^{-1}=\{0\}$ for any $\ell\geq 1$. Therefore, $I_{k+\ell}=\{0\}$ for any $\ell\geq 1$. This contradicts our hypothesis. Therefore, $\mathrm{deg}\,\Lambda/P<\mathrm{deg}\,\Lambda$ for any prime ideal $P$ containing $I_{k}$ for some $k$.
Let $n=\mathrm{deg}\,\Lambda$. Let $E(\Lambda)$ be the nonzero ideal generated by the evaluations in $\Lambda$ of $s_{2n-2}$ (cf. [MR, 13.6.7 Corollary]). We have just proved that for any fixed $k$, and for any prime ideal $P$ containing $I_{k}$, the ideal $E(\Lambda)$ is contained in $P$. Therefore, $E(\Lambda)$ is contained in the prime radical of $I_{k}$. Therefore, for any $k$ there exists $m_{k}$ such that $E(\Lambda)^{m_{k}}\subseteq I_{k}$ (recall that the prime radical of $\Lambda/I_{k}$ is nilpotent by [GW, Theorem 2.11]).
By hypothesis, $\Lambda/E(\Lambda)$ satisfies condition ()(*)( ∗ ). This implies that there exists $k_{0}$ such that $I_{k_{0}+\ell}+E(\Lambda)=I_{k_{0}}+E(\Lambda)$ for any $\ell\geq 0$. By Lemma 4.5, we deduce that $I_{k_{0}+\ell+1}$ is an idempotent ideal for any $\ell\geq 0$. As a consequence, the sequence (9) contains infinitely many different idempotent ideals, which gives the desired contradiction with the statement of [RS, Theorem 4]. This concludes the proof of the statement.
Corollary 4.9.
Let $\Lambda$ be a right noetherian PI-ring. Then:
  • (i)
    Every idempotent ideal of $\Lambda$ is the trace of a countably generated projective right $\Lambda$-module and of a countably generated projective left $\Lambda$-module.
  • (ii)
    $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\cong V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$.
Proof.  Statement $(i)$ follows from the fact that $\Lambda/N(\Lambda)$, where $N(\Lambda)$ denotes the prime radical of $\Lambda$, is two-sided noetherian [MR, 13.6.15], and that $N(\Lambda)$ is nilpotent. Therefore, if $I$ is an idempotent ideal of $\Lambda$, then $(I+N(\Lambda))/N(\Lambda)$ is an idempotent ideal of $\Lambda/N(\Lambda)$, so it is the trace of a countably generated projective right $\Lambda/N(\Lambda)$-module and the trace of a countably generated projective left $\Lambda/N(\Lambda)$-module (cf. [whitehead]). By Lemma 3.1, $I$ is also the trace ideal of a countably generated projective module on both sides.
By Theorem 4.8, $\Lambda$ and $\Lambda/N$ satisfy the ()(*)( ∗ )-condition and the symmetric ()(**)( ∗ ∗ )-condition. By Theorem 2.13, $V^{*}(\Lambda/N)=V(\Lambda/N)\sqcup B(\Lambda/N)$ and from Corollary 3.3 we deduce that $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}^{*}(\Lambda)=V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$.
Since $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$, the rest of the statement follows from Corollary 2.12.
It is an interesting open question to determine whether there exists a right noetherian ring $\Lambda$ with an idempotent ideal that is the trace of a (countably generated) projective, left $\Lambda$-module but not the trace of a projective right $\Lambda$-module. Corollary 4.9 shows that if such an example exists, it cannot be a PI-ring.
The following corollary was our initial motivation to develop these results. As it shows that the basics of the machinery developed in [wiegand] for the endomorphism ring of finitely generated modules over commutative noetherian rings of Krull dimension $1$, is also available in arbitrary Krull dimension.
Corollary 4.10.
Let $R$ be a commutative noetherian ring $R$, and let $\Lambda$ be a module-finite $R$-algebra. Then $\Lambda$ satisfies condition ()(*)( ∗ ) and its symmetric condition ()(**)( ∗ ∗ ). Moreover, it has only a finite number of idempotent ideals, and projective modules satisfy the conclusions of Corollary 4.9.
Proof.  If $\Lambda$ is generated by $n$-elements as $R$-module, then $\Lambda$ is a sub-quotient of $M_{n}(R)$, so it is PI and noetherian (cf. [MR, 13.4.9 Corollary]). Then, Theorem 4.8 implies that $\Lambda$ satisfies condition ()(*)( ∗ ).
The statement about idempotent ideals is a consequence of [RS, Theorem 4]
Even for a two-sided noetherian ring $\Lambda$, there may be a strict inclusion $V(\Lambda)\sqcup B(\Lambda)\subsetneq V^{*}(\Lambda)$. See, for example, the last two sections of [traces].
4.3. An application to direct sums of finitely generated modules
To transfer the information about projective modules over the endomorphism ring of a finitely generated module $M$ to $\mbox{\rm Add}(M)$, we shall use the following well-known result.
Proposition 4.11.
([libro, Theorem 4.7]) Let $R$ be a ring, and let $M$ be a right $R$-module with endomorphism ring $\Lambda=\mathrm{End}_{R}(M)$. Then the functor $\mathrm{Hom}_{R}(M,-)$ induces a category equivalence between $\mbox{\rm add}(M)$ and the category of finitely generated projective right $\Lambda$-modules.
Assume, in addition, that $M_{R}$ is finitely generated. Then the functor $\mathrm{Hom}_{R}(M,-)$ induces a category equivalence between $\mbox{\rm Add}(M)$ and the category of projective right $\Lambda$-modules.
Remark 4.12.
In the notation of Proposition 4.11, the inverse of the equivalence induced by $\mathrm{Hom}_{R}(M,-)$ is given by the functor $-\otimes_{\Lambda}M$.
Corollary 4.13.
Let $R$ be a commutative noetherian ring, and let $\Lambda$ be a module-finite $R$-algebra. Let $M$ be a finitely generated right $\Lambda$-module. Then any indecomposable element of $\mbox{\rm Add}(M)$ is finitely generated.
Proof.  As $\mathrm{End}_{\Lambda}(M)$ is a module-finite $R$-algebra, we can apply Corollary 4.10, to deduce that all countably generated projective modules over $\mathrm{End}_{\Lambda}(M)$ that are not finitely generated are relatively big. Hence, they cannot be indecomposable. The conclusion follows from Proposition 4.11.
Remark 4.14.
In general, if $\Lambda$ is a PI-ring, then the endomorphism ring of a finitely generated right $\Lambda$-module is also a PI-ring, see the argument of [PS, Theorem 2].
5. Projective versus locally projective modules
Let $R$ be a commutative domain with field of fractions $Q$. An $R$-module $M$ is torsion-free if the natural map $M\to M\otimes_{R}Q$ is injective.
Notation 5.1.
To ease the notation, we will usually write $M_{Q}$ for the localization of an $R$-module $M$ at $R\setminus\{0\}$.
In the following result, we recall that localization commutes with the $\mathrm{Hom}$ functor for suitable torsion-free modules. This property is crucial for the rest of the paper.
Lemma 5.2.
[AHP, Lemma 2.8] Let $R$ be a domain, and let $\Lambda$ be an $R$-algebra. Let $\Sigma$ be a multiplicative subset of $R$. If $M$ is a finitely generated right $\Lambda$-module that is torsion-free as an $R$-module, then the canonical injective homomorphism
$\varphi\colon\operatorname{Hom}_{\Lambda}(M,N)\otimes_{R}R_{\Sigma}\to\operatorname{Hom}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma},N\otimes_{R}R_{\Sigma})$
is an isomorphism for every $\Lambda$-module $N$ which is torsion-free as an $R$-module.
The following result is an extension of [wiegand, Lemma 10.1].
Lemma 5.3.
Let $R$ be a commutative ring, and let $\Lambda$ be an $R$-algebra. Let $R\to T$ be a flat ring homomorphism.
  1. (i)
    Let $P$ be a countably generated projective right $\Lambda$-module. Then $P\otimes_{R}T$ is a countably generated projective right module over $\Lambda\otimes_{R}T$ and
    $\operatorname{Tr}_{\Lambda\otimes_{R}T}(P\otimes_{R}T)=\operatorname{Tr}_{\Lambda}(P)\otimes_{R}T.$
  2. (ii)
    Let $T$ be a commutative ring. If $M$ is a finitely presented right $\Lambda$-module, then
    $\operatorname{Tr}_{\Lambda\otimes_{R}T}(M\otimes_{R}T)=\operatorname{Tr}_{\Lambda}(M)\otimes_{R}T.$
  3. (iii)
    Let $T=R_{\Sigma}$, where $\Sigma$ is a multiplicative subset of $R$ consisting of non-zero divisors. Assume $\Lambda$ is torsion-free as an $R$-module. If $M$ is a finitely generated right $\Lambda$-module which is torsion-free as an $R$-module, then
    $\operatorname{Tr}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma})=\operatorname{Tr}_{\Lambda}(M)\otimes_{R}R_{\Sigma}.$
Proof.  Statements $(i)$ and $(ii)$ are proved in [wiegand, Lemma 10.1].
To prove $(iii)$, let $x\in\operatorname{Tr}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma})$. Then there is an element $t\in\Sigma$, and $f_{1},\dotsc,f_{n}\in\operatorname{Hom}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma},\Lambda\otimes_{R}R_{\Sigma})$, $m_{1},\dotsc,m_{n}\in M$ such that $x=\sum_{i=1}^{n}f_{i}\left(m_{i}\otimes\frac{1}{t}\right)$. Let
$\varphi\colon\operatorname{Hom}_{\Lambda}(M,\Lambda)\otimes_{R}R_{\Sigma}\to\operatorname{Hom}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma},\Lambda\otimes_{R}R_{\Sigma})$
denote the canonical homomorphism. By Lemma 5.2, $\varphi$ is an isomorphism, so there is an element $r\in\Sigma$ and $g_{1},\dotsc,g_{n}\in\operatorname{Hom}_{\Lambda}(M,\Lambda)$ such that $\varphi(g_{i}\otimes\frac{1}{r})=f_{i}$, for each $i=1,\dotsc,n$. Then
$x=\sum_{i=1}^{n}f_{i}\left(m_{i}\otimes\frac{1}{t}\right)=\sum_{i=1}^{n}\varphi\left(g_{i}\otimes\frac{1}{r}\right)\left(m_{i}\otimes\frac{1}{t}\right)=\sum_{i=1}^{n}g_{i}(m_{i})\otimes\frac{1}{rt}.$
Therefore, $x\in\operatorname{Tr}_{\Lambda}(M)\otimes_{R}R_{\Sigma}$.
To prove the reverse containment, let $x\in\operatorname{Tr}_{\Lambda}(M)$ and $t\in R_{\Sigma}$. Then there are $f_{1},\dotsc,f_{n}\in\operatorname{Hom}_{\Lambda}(M,\Lambda)$, $m_{1},\dotsc,m_{n}\in M$ such that $x=\sum_{i=1}^{n}f_{i}(m_{i})$. Therefore, $x\otimes t=\sum_{i=1}^{n}f_{i}(m_{i})\otimes t=\sum_{i=1}^{n}\varphi(f_{i}\otimes 1)(m_{i}\otimes t)$. Hence, $x\otimes t\in\operatorname{Tr}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma})$. This concludes the proof of the statement.
The following result is necessary to give a version of the Package Deal Theorem for projective modules.
Corollary 5.4.
[AHP, Corollary 4.3] Let $R$ be a commutative domain, and let $\Lambda$ be an $R$-algebra. Let $N$ be a finitely generated right $\Lambda$-module which is torsion-free as an $R$-module. Then $N$ is a projective $\Lambda$-module if and only if $N_{\mathfrak{m}}$ is a projective $\Lambda_{\mathfrak{m}}$-module for every maximal ideal $\mathfrak{m}$ of $R$.
Definition 5.5.
A commutative domain $R$ is said to have finite character if any non-zero element is contained in only a finite number of maximal ideals. A commutative domain of finite character is said to be $h$-local if, in addition, any non-zero prime ideal of $R$ is contained in a unique maximal ideal.
The domain $R$ has finite character if $R/I$ is a semilocal ring for every non-zero ideal $I$ of $R$. If $R$ is $h$-local then it also satisfies that $R/\mathfrak{p}$ is a local domain for every non-zero prime ideal $\mathfrak{p}$ of $R$. Equivalently, $R$ is an $h$-local domain if and only if $R/I$ is semiperfect for every non-zero ideal $I$ of $R$. For characterizations of $h$-local domains the reader is referred to [olberding].
In general, one has the following easy property for modules over torsion-free algebras over a domain.
Lemma 5.6.
Let $R$ be a domain, and let $\Lambda$ be an $R$-algebra. Let $\Sigma$ be a multiplicative subset of $R$. Let $U$ be a right $\Lambda$-module which is torsion-free as an $R$-module. If $V$ is a finitely generated $\Lambda$-submodule that is torsion-free and such that $V_{\Sigma}$ is isomorphic to a submodule of $U_{\Sigma}$, then $U$ has a submodule isomorphic to $V$.
Proof.  By hypothesis, and because $V$ is torsion-free, $U_{\Sigma}$ has a submodule $V^{\prime}$ isomorphic to $V$. Since $V^{\prime}$ is finitely generated, there exists $d\in\Sigma$ such that $V^{\prime}\subseteq\frac{1}{d}U\subseteq U_{\Sigma}$. Then $dV^{\prime}\leq U$ is the module we are looking for.
If in Lemma 5.6, $R$ has finite character then the element $d$ that appears in the proof is contained only in a finite number of maximal ideals. So one easily proves results like the following one, which are crucial in our discussion.
Lemma 5.7.
[AHP, Lemma 4.2 and 5.1] Let $R$ be a commutative domain of finite character with field of fractions $Q$. Let $\Lambda$ be an $R$-algebra, and let $M,N$ be right $\Lambda$-modules which are torsion-free as $R$-modules. Assume that $M_{Q}\cong N_{Q}$. Then
  • (i)
    If $M$ and $N$ are finitely generated, then $M_{\mathfrak{m}}\cong N_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$.
  • (ii)
    If $M$ is finitely generated and $N\subseteq M$ then $M_{\mathfrak{m}}=N_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$. So that, there exist $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ maximal ideals of $R$ such that $M/N=\prod_{i=1}^{\ell}M_{\mathfrak{m}_{i}}/N_{\mathfrak{m}_{i}}$.
In the following result, we specialize to two-sided ideals of $\Lambda$ and assume that $\Lambda_{Q}$ is simple artinian.
Lemma 5.8.
Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian, and let $I$ be a non-zero two-sided ideal of $\Lambda$. Then:
  1. (i)
    $I\cap R\neq\{0\}$;
  2. (ii)
    $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$;
  3. (iii)
    there exists $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ maximal ideals of $R$ such that $\Lambda/I=\prod_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$.
Proof.  $(i)$ Since $\Lambda_{Q}$ is simple artinian and $\Lambda$ is torsion-free, $I_{Q}$ is a non-zero two-sided ideal of $\Lambda_{Q}$. So $I_{Q}=\Lambda_{Q}$. Therefore, there exist $x_{1},\dots,x_{n}\in I$ and $d\in R\setminus\{0\}$ such that $1=\sum_{i=1}^{n}\frac{x_{i}}{d}$. Thus $d\in R\cap I$ and this finishes the proof of $(i)$. Note that $d$ is contained only in a finite number of maximal ideals, so this yields $(ii)$.
Statement $(iii)$ is a direct application of Lemma 5.7.
The following results, which are known as Package Deal Theorems, were proved in [levyodenthal2] for the case of algebras over noetherian rings of Krull dimension 1. The name of these theorems comes from the idea of having a “package” of modules over a localization and finding a module over the non-localized algebra whose localizations are those given in the “package”. In [AHP] Package Deal Theorems were given for localizations of submodules and localizations of trace ideals. We recall here the statements.
Package Deal Theorem 5.9 (Localization of Submodules).
[AHP, Theorem 5.7] Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be an $R$-algebra, and let $M$ be a finitely generated right $\Lambda$-module, which is torsion-free as an $R$-module. For each maximal ideal $\mathfrak{m}$ of $R$, let $X(\mathfrak{m})$ be a $\Lambda_{\mathfrak{m}}$-submodule of $M_{\mathfrak{m}}$, which is torsion-free as an $R_{\mathfrak{m}}$-module, and such that $X(\mathfrak{m})_{Q}=(M_{\mathfrak{m}})_{Q}$. Then the following statements are equivalent
  1. (i)
    There is a $\Lambda$-submodule $N\subseteq M$, which is torsion-free as an $R$-module, and such that $N_{\mathfrak{m}}=X(\mathfrak{m})$ for all maximal ideals $\mathfrak{m}$ of $R$.
  2. (ii)
    $X(\mathfrak{m})=M_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$.
Moreover, if each of the $X(\mathfrak{m})$ is finitely generated as a $\Lambda_{\mathfrak{m}}$-module, then $N$ is also finitely generated as a $\Lambda$-module.
Package Deal Theorem 5.10 (Localization of Trace Ideals).
[AHP, Theorem 5.13] Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a torsion-free $R$-algebra such that $\Lambda_{Q}$ is a simple artinian ring. For each maximal ideal $\mathfrak{m}$ of $R$, let $I(\mathfrak{m})$ be a non-zero two-sided ideal of $\Lambda_{\mathfrak{m}}$ which is the trace ideal of a countably generated projective right $\Lambda_{\mathfrak{m}}$-module. Then the following statements are equivalent
  1. (i)
    There is a two-sided ideal $I$ of $\Lambda$, which is the trace ideal of a countably generated projective right $\Lambda$-module, and such that $I_{\mathfrak{m}}=I(\mathfrak{m})$ for all maximal ideals $\mathfrak{m}$ of $R$.
  2. (ii)
    $I(\mathfrak{m})=\Lambda_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$.
Moreover, the ideal $I$ that satisfies the equivalent conditions $(i)$ and $(ii)$ is unique.
Corollary 5.11.
Let $R$ be a commutative domain of finite character with field of fractions $Q$. Let $\Lambda$ be a torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $S$ denote a simple right $\Lambda_{Q}$-module, and assume that $\Lambda_{Q}\cong S^{k}$ as right $\Lambda_{Q}$-modules. Then any finitely generated projective right $\Lambda$-module $P$ satisfies that $P_{\mathfrak{m}}^{k}$ is free for almost all maximal ideals $\mathfrak{m}$ of $R$.
Proof.  By assumption $P_{Q}\cong S^{\ell}$ for suitable $\ell$. Therefore $P_{Q}^{k}\cong\Lambda_{Q}^{\ell}$. Now the conclusion follows from Lemma 5.7 using the fact that finitely generated projective modules over $\Lambda$ are torsion-free because $\Lambda$ is torsion-free.
Package Deal Theorem 5.12 (Localization of Projective Modules).
Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. For each maximal ideal $\mathfrak{m}$ of $R$, let $P(\mathfrak{m})$ be a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module with trace ideal $I(\mathfrak{m})$, and such that $P(\mathfrak{m})_{Q}\cong\Lambda_{Q}^{r}$ for some positive integer $r$. Then the following statements are equivalent:
  1. (i)
    There is a finitely generated projective right $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong P(\mathfrak{m})$ for every maximal ideal $\mathfrak{m}$ of $R$.
  2. (ii)
    $P(\mathfrak{m})\cong\Lambda_{\mathfrak{m}}^{r}$ for almost all maximal ideals $\mathfrak{m}$ of $R$.
Moreover, if $P$ is a finitely generated projective satisfying $(i)$ and $I=\operatorname{Tr}_{\Lambda}(P)$, then $I_{\mathfrak{m}}=I(\mathfrak{m})$ for every $\mathfrak{m}\in\operatorname{mSpec}\ R$.
Proof.  $(ii)\Rightarrow(i)$. Let $\mathcal{M}$ be the finite set of maximal ideals of $R$ such that $P(\mathfrak{m})\not\cong\Lambda_{\mathfrak{m}}^{r}$. Consider a module $F=\Lambda^{r}$. Since $\Lambda$ is torsion-free as an $R$-module, for every $\mathfrak{m}\in\mathcal{M}$ there exists a submodule $F(\mathfrak{m})$ of $F_{\mathfrak{m}}$ such that $F(\mathfrak{m})\cong P(\mathfrak{m})$. For $\mathfrak{n}\in\operatorname{mSpec}R\setminus\mathcal{M}$ set $F(\mathfrak{n})=F_{\mathfrak{n}}$. By Package Deal Theorem 5.9, there is a finitely generated right $\Lambda$-module $P$ contained in $F$ such that $P_{\mathfrak{m}}=F(\mathfrak{m})\cong P(\mathfrak{m})$ for every $\mathfrak{m}\in\mathcal{\operatorname{mSpec}}R$. By Corollary 5.4, $P$ is a projective module. Also note that $\operatorname{Tr}_{\Lambda}(P)_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}})$ by Lemma 5.3.
$(i)\Rightarrow(ii)$ Apply Lemma 5.7.
Corollary 5.13.
Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $S$ denote a simple right $\Lambda_{Q}$-module, and assume that $\Lambda_{Q}\cong S^{k}$ as right $\Lambda_{Q}$-modules. For each maximal ideal $\mathfrak{m}$ of $R$, let $P(\mathfrak{m})$ be a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module with trace ideal $I(\mathfrak{m})$. Assume there exists a positive integer $r$ such that $P(\mathfrak{m})_{Q}\cong S^{r}$. Then the following statements are equivalent:
  1. (i)
    There is a finitely generated projective right $\Lambda$-module $P$ with trace ideal $I$ such that $P_{\mathfrak{m}}\cong P(\mathfrak{m})^{k}$ and $I_{\mathfrak{m}}=I(\mathfrak{m})$ for every maximal ideal $\mathfrak{m}$ of $R$.
  2. (ii)
    $P(\mathfrak{m})^{k}$ is free for almost all maximal ideals $\mathfrak{m}$ of $R$.
Proof.  Apply Theorem 5.12 to a package $P(\mathfrak{m})^{\prime}:=P(\mathfrak{m})^{k}$, for each maximal ideal $\mathfrak{m}$ of $R$.
6. Semiperfect rings and locally semiperfect rings
We start this section by examining quickly how the monoids of projective modules we have introduced are for the case of semiperfect rings. We follow the notation and conventions of [andersonfuller, §27].
6.1. Projective modules over semiperfect rings and their trace ideals
Lemma 6.1.
Let $\Lambda$ be a semiperfect ring, and let $I$ be an ideal of $\Lambda$. Let $P^{\prime}$ be a finitely generated, projective right $\Lambda/I$-module. Then there exists a finitely generated projective right $\Lambda$-module $X$ such that $X/XI\cong P^{\prime}$.
Proof.  Since $\Lambda$ is semiperfect, $P^{\prime}$ has a projective cover. That is, it fits in an exact sequence
00{0}${\operatorname{Ker}f}$${X}$${P^{\prime}}$${0,}$$\scriptstyle{f}$
where $X$ is a finitely generated projective right $\Lambda$-module, and $\operatorname{Ker}f$ is a small submodule of $X$. Therefore, there are inclusions
$XI\subseteq\operatorname{Ker}f\subseteq XJ(\Lambda)$
which induce an exact sequence of $\Lambda/I$-modules
00{0}${\operatorname{Ker}f/XI}$${X/XI}$${P^{\prime}}$${0.}$$\scriptstyle{f^{\prime}}$
which splits because $P^{\prime}$ is projective. Since $\operatorname{Ker}f/XI$ is also a small submodule of $X/XI$, it must be zero. Therefore $X/XI\cong P^{\prime}$.
In the next proposition, we recall the behavior of traces of projective modules for semiperfect rings.
Proposition 6.2.
Let $\Lambda$ be a semiperfect ring, and let $J(\Lambda)$ denote its Jacobson radical. Then:
  1. (i)
    $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda):=\mathcal{T}(\Lambda)$, and the injective map $\varphi\colon\mathcal{T}(\Lambda)\to\mathcal{T}(\Lambda/J(\Lambda))$ defined in Lemma 1.8 is bijective.
  2. (ii)
    Let $e_{1},\dots,e_{t}$ be a basic set of (primitive orthogonal) idempotents of $\Lambda$. Let $\mathcal{P}(\{1,\dots,t\})$ denote the set of all subsets of $\{1,\dots,t\}$. Then the map
    $\Gamma\colon\mathcal{P}(\{1,\dots,t\})\to\mathcal{T}(\Lambda)$
    defined by $\Gamma(A)=\sum_{i\in A}\Lambda e_{i}\Lambda$, for any $A\in\mathcal{P}(\{1,\dots,t\})$, is a bijection.
  3. (iii)
    Let $P$ be a projective right $\Lambda$-module with trace $I$. Assume that $I/IJ(\Lambda)\cong\sum_{i\in A}\left((e_{i}+J(\Lambda))\Lambda/J(\Lambda)\right)^{\ell_{i}}$ for suitable $\ell_{i}>0$ and $A\in\mathcal{P}(\{1,\dots,t\})$. Then $P\cong\bigoplus_{i\in A}\left(e_{i}\Lambda\right)^{(L_{i})}$, for suitable non-empty sets, $L_{i}$.
  4. (iv)
    Keeping the notation from $(ii)$ and $(iii)$, $\Lambda^{\prime}:=\Lambda/I$ is a semiperfect ring and $e_{i}^{\prime}:=e_{i}+I,i\not\in A$ is a basic set of idempotents of $\Lambda/I$. Moreover, let $P^{\prime}\cong\bigoplus_{i\not\in A}(e_{i}^{\prime}\Lambda^{\prime})^{x_{i}}$ be a finitely generated projective $\Lambda^{\prime}$-module and let $P$ be a finitely generated projective $\Lambda$-module such that $P/PI\cong P^{\prime}$. Then there are non-negative integers $s_{i},i\not\in A$ such that $P\cong\bigoplus_{i\not\in A}(e_{i}\Lambda)^{x_{i}}\oplus\bigoplus_{i\in A}(e_{i}\Lambda)^{s_{i}}$.
Proof.  The proof of the four statements is more or less routine. One just has to take into account that any projective right (left) $\Lambda$-module is isomorphic to a unique direct sum of the indecomposable projective modules $P_{i}=e_{i}\Lambda$ ($Q_{i}=\Lambda e_{i}$), for $i=1,\dots,t$ (see for example [libro, Theorem 3.10]). Moreover $\mathrm{Tr}_{\Lambda}(P_{i})=\mathrm{Tr}_{\Lambda}(Q_{i})=\Lambda e_{i}\Lambda$.
Proposition 6.3.
Let $\Lambda$ be a semiperfect ring. Then:
  • (i)
    The canonical isomorphism between $V_{r}(\Lambda):=V(\Lambda)\cong{\mathbb{N}}_{0}^{t}$ and $V_{\ell}(\Lambda)$ extends to an isomorphism $V_{r}^{*}(\Lambda)\cong V_{\ell}^{*}(\Lambda):=V^{*}(\Lambda)\cong({\mathbb{N}}_{0}^{*})^{t}$;
  • (ii)
    Let $P$ be a projective right $\Lambda$-module. Then there exists a two-sided ideal $I$ of $\Lambda$ minimal with respect to the property $P/PI$ is finitely generated. Moreover, $I$ is the trace of a finitely generated projective module $Q$ and $P$ is $I$-big.
  • (iii)
    $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$.
Proof.  $(i)$ Let $e_{1},\dots,e_{t}$ be a basic set of idempotents of $\Lambda$. Then $P_{i}=e_{i}\Lambda$ and $Q_{i}=\Lambda e_{i}\cong\mathrm{Hom}_{\Lambda}(P_{i},\Lambda)$, for $i=1,\dots,t$, are representatives of the isomorphism classes of indecomposable projective right and left $\Lambda$-modules. As any projective module is a direct sum of indecomposables, in a unique way up to isomorphism, the first statement is clear.
$(ii)$ If $P$ is a projective right $\Lambda$-module, then $\{1,\dots,t\}=A\cup B$, where $A$ and $B$ are disjoint, and $P\cong\left(\bigoplus_{i\in A}P_{i}^{(A_{i})}\right)\oplus\left(\bigoplus_{i\in B}P_{i}^{(B_{i})}\right)$, where $A_{i}$ is a finite set (possibly empty) for any $i\in A$ and $B_{i}$ is an infinite set for any $i\in B$. Then
$I=\sum_{i\in B}\mathrm{Tr}_{\Lambda}(P_{i})=\mathrm{Tr}_{\Lambda}\Big{(}\bigoplus_{i\in B}P_{i}\Big{)}$
is the ideal we are looking for.
$(iii)$ By Proposition 6.2, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ so we can apply Corollary 2.12 to deduce that $B_{r}(\Lambda):=B(\Lambda)$ is canonically isomorphic to $B_{\ell}(\Lambda)$. The statement $(ii)$ implies that $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$.
6.2. Locally semiperfect algebras over $h$-local domains
Definition 6.4.
Let $\Lambda$ be an algebra over a commutative ring $R$. We say that $\Lambda$ is locally semiperfect if $\Lambda_{\mathfrak{m}}$ is semiperfect for every maximal ideal $\mathfrak{m}$ of $R$.
In this section, it is important to keep in mind that we are using the conventions of Notation 5.1.
The following example explains a particular instance of the appearance of locally semiperfect algebras in which we will be particularly interested in Section 7.
Example 6.5.
Let $R$ be a commutative domain with field of fractions $Q$, and let $M$ be a finitely generated torsion-free $R$-module. Then $\Lambda=\operatorname{End}_{R}(M)$ is an $R$-algebra and, by Lemma 5.2, $\Lambda_{Q}=\Lambda\otimes_{R}Q\cong\operatorname{End}_{Q}(M\otimes_{R}Q)=\operatorname{End}_{Q}(M_{Q})$ is simple artinian. By [libro, Proposition 3.14] and Lemma 5.2, for every maximal ideal $\mathfrak{m}$ of $R$, $M_{\mathfrak{m}}$ is a direct sum of modules with local endomorphism ring if and only if $\Lambda_{\mathfrak{m}}\cong\operatorname{End}_{R_{\mathfrak{m}}}(M_{\mathfrak{m}})$ is a semiperfect ring. In other words, $\Lambda$ is locally semiperfect if and only if $M_{\mathfrak{m}}$ is a direct sum of modules with local endomorphism ring for every maximal ideal $\mathfrak{m}$ of $R$.
Lemma 6.6.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Then every trace of a projective right (or left) $\Lambda$-module is also the trace of a finitely generated projective right $\Lambda$-module. In particular, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$.
Proof.  Let $I$ be a non-zero two-sided ideal of $\Lambda$ which is the trace of a projective right $\Lambda$-module. Since $\Lambda$ is torsion-free, $I_{\mathfrak{m}}\neq\{0\}$ for every maximal ideal $\mathfrak{m}$ of $R$. By Lemma 5.3 and Lemma 6.2, $I_{\mathfrak{m}}$ is also the trace of a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module.
Let $S$ denote a simple right $\Lambda_{Q}$-module, and assume that $\Lambda_{Q}\cong S^{k}$ as right $\Lambda_{Q}$-modules. Since $\Lambda_{Q}$ is simple artinian and $I$ is non-zero, by Lemma 5.8, there exists a finite set $\mathcal{M}$ of maximal ideals of $R$ such that $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ if $\mathfrak{m}\not\in\mathcal{M}$. For any maximal ideal $\mathfrak{m}\in\mathcal{M}$, let $P(\mathfrak{m})$ be a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module with trace ideal $I_{\mathfrak{m}}$. Then $P(\mathfrak{m})_{Q}\cong S^{t_{\mathfrak{m}}}$ for a suitable positive integer $t_{\mathfrak{m}}\geq 1$. Let $t=\mathrm{lcm}\{t_{\mathfrak{m}}\mid\mathfrak{m}\in\mathcal{M}\}$, write $t=\ell_{\mathfrak{m}}\cdot t_{\mathfrak{m}}$, and let $P^{\prime}(\mathfrak{m}):=P(\mathfrak{m})^{t\ell_{\mathfrak{m}}}$ for each maximal ideal $\mathfrak{m}$ of $\mathcal{M}$. For each $\mathfrak{m}\notin\mathcal{M}$, take $P^{\prime}(\mathfrak{m})=\Lambda_{\mathfrak{m}}^{t}$. Note that $P^{\prime}(\mathfrak{m})_{Q}\cong S^{tk}\cong\Lambda_{Q}^{t}$, and $\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P^{\prime}(\mathfrak{m}))=I_{\mathfrak{m}}$ for every maximal ideal $\mathfrak{m}$ of $R$. By Package Deal Theorem 5.12 applied the package $\{P^{\prime}(\mathfrak{m})\mid\mathfrak{m}\in\operatorname{mSpec}R\}$ there exists a finitely generated projective right $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong P^{\prime}(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. By Lemma 5.3, $\operatorname{Tr}_{\Lambda}(P)_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P^{\prime}(\mathfrak{m}))=I_{\mathfrak{m}}$ for any maximal ideal $\mathfrak{m}$ of $R$. This implies that $I=\operatorname{Tr}_{\Lambda}(P)+I=\operatorname{Tr}_{\Lambda}(P)$.
A similar argument gives the conclusion for the traces of projective left $\Lambda$-modules because of the symmetry of the hypotheses.
The equality $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ follows because of the canonical bijection between finitely generated projective right and left $\Lambda$-modules induced by the $\Lambda$-dual, and because the trace of a finitely generated projective module coincides with the trace of its $\Lambda$-dual, cf. Remark 2.9.
Corollary 6.7.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. If $P$ is a countably generated projective right $\Lambda$-module, then $P^{(\omega)}$ is a direct sum of finitely generated modules.
Proof.  By Lemma 6.6, the trace of $P$ coincides with the trace of a finitely generated projective module. To conclude, apply that if $P$ and $Q$ are countably generated projective modules then $P^{(\omega)}\cong Q^{(\omega)}$ if and only if $P$ and $Q$ have the same trace.
Lemma 6.8.
Let $R$ be a commutative domain of finite character with field of fractions $Q$. Let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian; so that there exist $k$ and a division ring D such that $\Lambda_{Q}\cong M_{k}(D)$. Then for almost all maximal ideals $\mathfrak{m}$ of $R$ the ring $\Lambda_{\mathfrak{m}}$ has the following properties:
  1. (i)
    $\Lambda_{\mathfrak{m}}$ has exactly one indecomposable projective up to an isomorphism. Moreover, $\Lambda_{\mathfrak{m}}$ is isomorphic to a ring of $k\times k$ matrices over a local ring $T_{\mathfrak{m}}$ and $T_{\mathfrak{m}}\otimes_{R}Q\cong D$.
  2. (ii)
    If $U$ and $V$ are projective $\Lambda_{\mathfrak{m}}$-modules then $U\cong V$ if and only if $U_{Q}\cong V_{Q}$.
Proof.  Let $S$ be a simple $\Lambda_{Q}$-module and assume that $\Lambda_{Q}\cong S^{k}$. Let $X$ be a finitely generated $\Lambda$-module, which is torsion-free as an $R$-module, and such that $X_{Q}\cong S$. By Lemma 5.7, $X_{\mathfrak{m}}^{k}$ is free for almost all maximal ideals. If $\mathfrak{m}$ is such a maximal ideal, then $X_{\mathfrak{m}}^{k}\cong\Lambda_{\mathfrak{m}}$ gives an indecomposable decomposition of the semiperfect ring $\Lambda_{\mathfrak{m}}$. Hence $X_{\mathfrak{m}}$ is the only indecomposable projective $\Lambda_{\mathfrak{m}}$-module up to an isomorphism. As $\Lambda_{\mathfrak{m}}$ is semiperfect, the endomorphism ring of $X_{\mathfrak{m}}$ is a local ring that we will denote by $T_{\mathfrak{m}}$. Hence $\Lambda_{\mathfrak{m}}\cong\mathrm{End}_{\Lambda_{\mathfrak{m}}}(X_{\mathfrak{m}}^{k})\cong M_{k}(T_{\mathfrak{m}}).$ By Lemma 5.2, $T_{\mathfrak{m}}\otimes_{R}Q\cong D$.
Projective modules over $M_{k}(T_{\mathfrak{m}})$ are always a direct sum of copies of the indecomposable projective $X_{\mathfrak{m}}$. Moreover, $U\cong X_{\mathfrak{m}}^{(I)},V\cong X_{\mathfrak{m}}^{(J)}$ then $U\cong V$ if and only if $I$ and $J$ have the same cardinality, if and only if $U_{Q}\cong V_{Q}$.
Remark 6.9.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. In view of Lemma 6.8, the maximum spectrum of $R$ can be written as a disjoint union of two sets $A$ and $B$ such that for any $\mathfrak{m}\in A$, $\Lambda_{\mathfrak{m}}$ is a ring of matrices over a local ring and $B$ is a finite, non-empty set. In particular, for $\mathfrak{m}\in A$, $\mathcal{T}_{r}(\Lambda_{\mathfrak{m}})=\mathcal{T}_{\ell}(\Lambda_{\mathfrak{m}})=\{0,\Lambda_{\mathfrak{m}}\}$.
In view of the Package Deal Theorem for Traces 5.10, there is a bijective correspondence
${\mathcal{T}_{r}(\Lambda)\setminus\{0\}=\mathcal{T}_{\ell}(\Lambda)\setminus\{0\}}$${\prod_{\mathfrak{m}\in B}\left(\mathcal{T}_{r}(\Lambda_{\mathfrak{m}})\setminus\{0\}\right).}$
Corollary 6.10.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Then $\mathcal{T}(\Lambda)$ is finite. Moreover, if $P$ is a finitely generated projective $\Lambda$-module, then $I=\operatorname{Tr}_{\Lambda}(P)$ is a minimal nonzero element of $\mathcal{T}(\Lambda)$ if and only if $P_{\mathfrak{m}}$ is a finite power of an indecomposable projective $\Lambda_{\mathfrak{m}}$-module for any $\mathfrak{m}\in\operatorname{mSpec}R$.
Proof.  Finiteness of ${\mathcal{T}}(\Lambda)$ follows from Remark 6.9. Also note that the correspondence of the remark preserves inclusions, so $\operatorname{Tr}_{\Lambda}(P)$ is a minimal element of ${\mathcal{T}}(\Lambda)\setminus\{0\}$ if and only if $\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}})$ is a minimal element of ${\mathcal{T}}(\Lambda_{\mathfrak{m}})\setminus\{0\}$ for any $\mathfrak{m}\in\operatorname{mSpec}(R)$ if and only if $P_{\mathfrak{m}}$ is a finite power of an indecomposable projective $\Lambda_{\mathfrak{m}}$-module.
Proposition 6.11.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $P$ be a countably generated projective right $\Lambda$-module. Then,
  1. (1)
    $P_{\Lambda}$ is finitely generated if and only if $P_{Q}$ is a finitely generated $\Lambda_{Q}$-module if and only if there exists a maximal ideal $\mathfrak{m}$ of $R$ such that $P_{\mathfrak{m}}$ is a finitely generated $\Lambda_{\mathfrak{m}}$-module.
  2. (2)
    The set $\mathcal{C}=\{J\unlhd\Lambda\mid P/PJ\mbox{ is finitely generated}\}$ has a minimal element which is the trace of a finitely generated, projective right $\Lambda$-module. If $I$ is such ideal then it is characterized by the property that, for any maximal ideal $\mathfrak{m}$ of $R$, $I_{\mathfrak{m}}$ is the minimal element of $\mathcal{C}_{\mathfrak{m}}=\{J\unlhd\Lambda\mid P_{\mathfrak{m}}/P_{\mathfrak{m}}J\mbox{ is finitely generated}\}$.
Proof.  Throughout the proof, let $S$ denote a simple right $\Lambda_{Q}$-module.
$(1)$ We claim that if $P_{Q}\cong S^{n}$ for some $n\in\mathbb{N}$, then $P$ is finitely generated. Indeed, by hypothesis, there exists a finitely generated $\Lambda$-submodule $P_{0}$ of $P$ such that $(P_{0})_{Q}=P_{Q}$. Then we may assume that we have, $P_{0}\leq P\leq\Lambda^{(\omega)}=P\oplus P^{\prime}$, and for any $p\in P$ there exists $0\neq d\in R$ such that $pd\in P_{0}$. Since $\Lambda$ is torsion-free as an $R$-module, this implies that there exists a finite subset $F$ of $\omega$ such that $P_{0}\leq P\subseteq\Lambda^{F}$. Hence $\Lambda^{F}=P\oplus(P^{\prime}\cap\Lambda^{F})$ so $P$ is finitely generated.
This proves statement $(1)$.
$(2)$ In the proof, we keep the notation of Remark 6.9 of the partition of the maximal spectrum of $R$.
If $P$ is finitely generated, the statement is clear. Assume that $P$ is a countably generated, projective, right $\Lambda$-module that is not finitely generated. By $(1)$, $P_{Q}\cong S^{(\omega)}$. Therefore, there are two possible situations with the localizations of $P$ at maximal ideals of $R$:
  • (i)
    $P_{\mathfrak{m}}$ is a free $\Lambda_{\mathfrak{m}}$-module of infinite rank for any $\mathfrak{m}\in A$; so $\Lambda_{\mathfrak{m}}$ is the only element of the set $\mathcal{C}_{\mathfrak{m}}=\{J\mbox{ ideal of $\Lambda_{\mathfrak{m}}$}\mid P_{\mathfrak{m}}/P_{\mathfrak{m}}J\mbox{ is finitely generated}\}$.
  • (ii)
    $P_{\mathfrak{n}}$ is an infinitely generated projective module over the semiperfect ring $\Lambda_{\mathfrak{n}}$ for any $\mathfrak{n}\in B$; by Proposition 6.3, there exists an ideal $I(\mathfrak{n})\neq\{0\}$, minimal element of the set $\mathcal{C}_{\mathfrak{n}}=\{J\mbox{ ideal of $\Lambda_{\mathfrak{n}}$}\mid P_{\mathfrak{n}}/P_{\mathfrak{n}}J\mbox{ is finitely generated}\}$. In addition $I(\mathfrak{n})$ is the trace of a finitely generated projective right $\Lambda_{\mathfrak{n}}$-module.
By the Package Deal Theorem 5.10, there exists $I$ ideal of $\Lambda$ such that $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ for any $\mathfrak{m}\in A$ and $I_{\mathfrak{n}}=I(\mathfrak{n})$ for any $\mathfrak{n}\in B$. In addition, $I$ is the trace of a (finitely generated, by Lemma 6.6) projective right $\Lambda$-module. We claim that $I$ is the minimal element of $\mathcal{C}$.
Since $I_{Q}=\Lambda_{Q}$, $I\cap R\neq\{0\}$. As $R$ has finite character, there exists a finite set $\mathcal{M}$ of maximal ideals of $R$, such that $R/(I\cap R)\cong\prod_{\mathfrak{m}\in\mathcal{M}}\left(R/(I\cap R)\right)_{\mathfrak{m}}$. We may assume that $B\subseteq\mathcal{M}$. Hence,
$P/PI\cong\prod_{\mathfrak{m}\in\mathcal{M}}P_{\mathfrak{m}}/P_{\mathfrak{m}}I_{\mathfrak{m}}=\prod_{\mathfrak{n}\in B}P_{\mathfrak{n}}/P_{\mathfrak{n}}I(\mathfrak{n})$
which is finitely generated. Therefore, $I\in\mathcal{C}$.
Let $I^{\prime}\in\mathcal{C}$ be such that $I^{\prime}\subseteq I$. Then, because $\Lambda$ is torsion-free as an $R$-module, for any maximal ideal $\mathfrak{m}$ of $R$, $I^{\prime}_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}$ and $(I^{\prime})_{\mathfrak{m}}\subseteq I_{\mathfrak{m}}=I(\mathfrak{m})$. The minimality in the choice of $I(\mathfrak{m})$ in $(i)$ and $(ii)$, implies that $(I^{\prime})_{\mathfrak{m}}=I_{\mathfrak{m}}=I(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. Hence, $I^{\prime}=I$. This shows that $I$ is the minimal element of $\mathcal{C}$.
Notice that the proof also shows that, for any maximal ideal $\mathfrak{m}$ of $R$, $I_{\mathfrak{m}}$ is the minimal element of $\mathcal{C}_{\mathfrak{m}}$, which is in $\mathcal{T}(\Lambda_{\mathfrak{m}})$ Proposition 6.3. By the Package Deal Theorem 5.10, $I$ is uniquely determined by its localizations.
Lemma 6.12.
Let $R$ be a commutative domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra. Let $P$ be a countably generated projective $\Lambda$-module, $I\in\mathcal{T}(\Lambda)$, and $V$ a finitely generated $\Lambda$-module which is torsion-free as an $R$-module. Let $\mathfrak{m}$ be a prime ideal of $R$ such that $V_{\mathfrak{m}}=V_{\mathfrak{m}}I_{\mathfrak{m}}$ and $P_{\mathfrak{m}}=\bigoplus_{i\in\mathbb{N}}U_{i}$ with $U_{i}$ finitely generated $\Lambda_{\mathfrak{m}}$-modules such that infinitely many of these $U_{i}$ are isomorphic to a $\Lambda_{\mathfrak{m}}$-module $U$ whose trace ideal contains $I_{\mathfrak{m}}$. Then for every finitely generated submodule $P_{1}$ of $P$, there exists a homomorphism $f\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $f(P_{1})=0$ and $f_{\mathfrak{m}}$ is onto.
Proof.  Let $\nu\in\operatorname{Hom}_{\Lambda}(P,\Lambda^{(\omega)})$ and $\pi\in\operatorname{Hom}_{\Lambda}(\Lambda^{(\omega)},P)$ be such that $\pi\nu=1_{P}$.
Note that if $U$ is a projective $\Lambda_{\mathfrak{m}}$-module and $I_{\mathfrak{m}}\subseteq\operatorname{Tr}_{\Lambda}(U)$, then $U$ generates any $\Lambda_{\mathfrak{m}}$-module $W$ satisfying $WI_{\mathfrak{m}}=W$. If $W$ is also finitely generated, then there exists $k\in\mathbb{N}$ such that $W$ is a homomorphic image of $U^{k}$. Therefore $V_{\mathfrak{m}}$ is a homomorphic image of $\bigoplus_{i\in C}U_{i}$ whenever $C$ is a cofinal subset of $\mathbb{N}$. In particular, there exists a homomorphism $g\in\operatorname{Hom}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}},V_{\mathfrak{m}})$ which is onto and $g(P_{1})=0$ (recall $P$ is embedded in $P_{\mathfrak{m}}$). Since $V$ is finitely generated, there exists $P_{0}^{\prime}$, a finitely generated submodule of $P_{\mathfrak{m}}$, such that $g(P_{0}^{\prime})=V_{\mathfrak{m}}$. Scaling by a suitable element of $R\setminus\mathfrak{m}$ (which is invertible over $\Lambda_{\mathfrak{m}}$) we observe that there exists finitely generated $P_{0}\subseteq P$ such that $P_{0}\Lambda_{\mathfrak{m}}=P_{0}^{\prime}$. Let $F$ be a finite subset of $\omega$ such that $\nu(P_{0}+P_{1})$ is contained in $\Lambda^{(F)}$, further let $\iota\in\operatorname{Hom}_{\Lambda}(\Lambda^{(F)},\Lambda^{(\omega)})$ and $\varrho\in\operatorname{Hom}_{\Lambda}(\Lambda^{(\omega)},\Lambda^{(F)})$ be the canonical embedding and projection. Since $\Lambda^{(F)}$ is finitely generated over $\Lambda$ and torsion-free over $R$, by Lemma 5.2 there exists $h\in\operatorname{Hom}_{\Lambda}(\Lambda^{(F)},V)$ and $q\in R_{\mathfrak{m}}$ such that $h\otimes q=g(\pi\iota)_{\mathfrak{m}}$. Consider $f:=h\varrho\nu$. Then $f\otimes q=h\otimes q\circ(\varrho\nu)\otimes 1=g(\pi\iota\varrho\nu)_{\mathfrak{m}}$. Since $\Lambda^{(F)}$ contains $\nu(P_{0}+P_{1})$, $\pi\iota\varrho\nu$ is identity on $P_{0}+P_{1}$ and, consequently, $f\otimes q$ is onto and $f\otimes q(P_{1}\otimes 1)=0$. This implies that $f_{\mathfrak{m}}$ is onto and $f(P_{1})=0$.
Remark 6.13.
We are going to apply Lemma 6.12 in the following situation: $\Lambda$ is a locally semiperfect torsion-free algebra over an $h$-local domain $R$ such that $\Lambda_{Q}$ is a simple artinian ring. Assume that $P$ is a countably generated projective $\Lambda$-module and $0\neq I$ is the least element of the set of ideals $\{J\unlhd\Lambda\mid P/PJ\text{ is finitely generated}\}$. Take any $\mathfrak{m}\in\operatorname{mSpec}(R)$ and let us check that $P_{\mathfrak{m}}$ is a countable direct sum of finitely generated projective $\Lambda_{\mathfrak{m}}$-modules such that almost all these direct summands have trace equal to $I_{\mathfrak{m}}$.
Indeed, since $P$ is not finitely generated, $P_{\mathfrak{m}}$ is not a finitely generated $\Lambda_{\mathfrak{m}}$-module by Proposition 6.11. Since $\Lambda_{\mathfrak{m}}$ is semiperfect, $P_{\mathfrak{m}}$ is a direct sum of finitely generated modules, $P_{\mathfrak{m}}=U_{1}^{m_{1}}\oplus\cdots\oplus U_{t}^{m_{t}}\oplus(U_{t+1}\oplus\cdots\oplus U_{s})^{(\omega)}$, where $U_{1},\dots,U_{s}$ is a representative set of indecomposable projective $\Lambda_{\mathfrak{m}}$-modules. Let $U^{\prime}=U_{t+1}\oplus\cdots\oplus U_{s}$. As explained in the proof of Proposition 6.11(2), $I_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(U^{\prime})$. Therefore $P_{\mathfrak{m}}=U_{1}^{m_{1}}\oplus\cdots\oplus U_{t}^{m_{t}}\oplus U^{\prime(\omega)}$ is a decomposition of $P_{\mathfrak{m}}$ having the required properties.
Lemma 6.12 can be stated in simpler form when $\mathfrak{m}=(0)$ and $I=\Lambda$.
Lemma 6.14.
Let $R$ be a commutative domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra. Let $P$ be a projective $\Lambda$-module such that $P_{Q}$ is countably generated but not finitely generated free, and $V$ a finitely generated $\Lambda$-module torsion-free over $R$. Then for any $P_{1}$ finitely generated submodule of $P$, there exists $f\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $f_{Q}$ is onto and $f(P_{1})=0$.
Proposition 6.15.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $P$ be a countably generated projective $\Lambda$-module such that $I=\min\{J\unlhd\Lambda\mid P/PJ\text{ is finitely generated}\}$ exists and it is the trace of a countably generated projective module. Then $P$ is $I$-big.
Proof.  We have to prove that if $U$ is a countably generated projective right $\Lambda$-module and $\operatorname{Tr}_{\Lambda}(U)\subseteq I$, then $U$ is a direct summand of $P$. By Corollary 6.7, $U^{(\omega)}$ is a direct sum of finitely generated modules, so we may assume that $U$ is a direct sum of finitely generated modules.
We prove the following claim: If $P_{0}$ is a finitely generated submodule of $P$ and $V$ is a finitely generated projective $\Lambda$-module with trace contained in $I$, then there exists an epimorphism $f\colon P\to V$ such that $f(P_{0})=0$.
Assume we have proved the claim and let $U=\bigoplus_{i\in\mathbb{N}}U_{i}$ with $U_{i}$ finitely generated. Let $p_{1},p_{2},\dots\in P$ be a countable set of generators of $P$. Consider the following construction: Let $f_{1}\colon P\to U_{1}$ be an epimorphism and let $P_{1}$ be a finitely generated submodule of $P$ such that $f_{1}(P_{1})=U_{1}$ and $p_{1}\in P_{1}$. Assume that $f_{1},f_{2},\dots,f_{k}$ and $P_{1},P_{2},\dots,P_{k}$ were defined such that $f_{i}\colon P\to U_{i}$ is an epimorphism, $P_{1}\subseteq\cdots\subseteq P_{k}$ are finitely generated submodules of $P$ such that $f_{i}(P_{i-1})=0$, $p_{i}\in P_{i}$, and $f_{i}(P_{i})=U_{i}$ (we can set $P_{0}=0$). By the claim, there exists an epimorphism $f_{k+1}\colon P\to U_{k+1}$ such that $f_{k+1}(P_{k})=0$. Finally, let $P_{k+1}$ be a finitely generated submodule of $P$ satisfying $P_{k}\subseteq P_{k+1}$, $p_{k+1}\in P_{k+1}$ and $f_{k+1}(P_{k+1})=U_{k+1}$.
Consider the canonical homomorphism $f=\prod_{k\in\mathbb{N}}f_{k}\colon P\to\prod_{k}U_{k}$. Observe that $f_{k}(p_{l})=0$ if $k>l$ so ${\rm Im}\ f\subseteq\bigoplus_{i\in\mathbb{N}}U_{i}$. An easy inductive argument also gives $f(P_{i})=\bigoplus_{j=1}^{i}U_{j}$ for any $i\in\mathbb{N}$. Thus $f$ is an epimorphism.
Thus we are left to prove the claim. First note that we may assume $I\neq 0$ since any countably generated projective module is 00-big. Then $P$ is not finitely generated and therefore, by Proposition 6.11, $P_{Q}$ cannot be finitely generated. Hence $P_{Q}\simeq\Lambda_{Q}^{(\omega)}$.
By Lemma 6.14, there exists $g\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $g_{Q}$ is onto and $g(P_{0})=0$. Let $T_{0}$ be a finitely generated submodule of $P$ containing $P_{0}$ such that $g_{Q}(T_{0}\Lambda_{Q})=(g(T_{0}))_{Q}=V_{Q}$. Since $V$ is finitely generated, $V_{\mathfrak{m}}=g(T_{0})_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$. Let $\mathfrak{m}_{1},\mathfrak{m}_{2},\dots,\mathfrak{m}_{k}$ be the maximal ideals of $R$ where the equality does not hold. If there are no such ideals then we are done, since $g$ is onto. So we may assume that $k\geq 1$.
It remains to show that there are homomorphisms $f_{1},f_{2},\dots,f_{k}\in\operatorname{Hom}_{\Lambda}(P,V)$ and finitely generated submodules $T_{1},T_{2},\dots,T_{k}$ of $P$ such that
  1. (i)
    $T_{0}\subseteq T_{1}\subseteq T_{2}\subseteq\cdots\subseteq T_{k}$
  2. (ii)
    $f_{i}(T_{i-1})=0$ for $i=1,\dots,k$
  3. (iii)
    $(g+\sum_{j=1}^{i}f_{j})_{\mathfrak{m}_{i}}((T_{i})_{\mathfrak{m}_{i}})=V_{\mathfrak{m}_{i}}$ for $i=1,\dots,k$
Then $f=g+\sum_{i=1}^{k}f_{i}\in\operatorname{Hom}_{\Lambda}(P,V)$ satisfies $f(T_{0})=g(T_{0})$ hence $f_{\mathfrak{m}}$ is onto if $\mathfrak{m}$ is not among $\mathfrak{m}_{1},\dots,\mathfrak{m}_{k}$. Moreover, $f(T_{i})=(g+\sum_{j=1}^{i}f_{j})(T_{i})$ shows that $f_{\mathfrak{m}_{i}}$ is onto for every $i=1,\dots,k$. This implies that $f$ is onto.
Assume $f_{1},\dots,f_{\ell-1}$ and $T_{0},T_{1},\dots,T_{\ell-1}$ are found. Consider $g_{\ell}=g+\sum_{i=1}^{\ell-1}f_{i}$. Recall that $V_{\mathfrak{m}_{\ell}}$ is a finitely generated projective module over a semiperfect ring $\Lambda_{\mathfrak{m}_{\ell}}$. Then $V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ is a finitely generated (semisimple module), so if $\pi\colon V_{\mathfrak{m}_{\ell}}\to V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ is the canonical projection, $\pi((g_{\ell}(P))_{\mathfrak{m}_{\ell}})$ is a finitely generated submodule of $V_{\mathfrak{m}_{\ell}}/V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$. Let $T_{\ell}^{\prime}\subseteq P$ be a finitely generated submodule of $P$ containing $T_{\ell-1}$ such that $(g_{\ell}(T_{\ell}^{\prime}))_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})=(g_{\ell}(P))_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$.
By Lemma 6.12, there exists $f_{\ell}\in{\rm Hom}_{\Lambda}(P,V)$ such that $(f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto and $f_{\ell}(T_{\ell}^{\prime})=0$. Then the image of $g_{\ell}+f_{\ell}$ contains $g_{\ell}(T_{\ell}^{\prime})$ hence also $\mbox{\rm Im}\ (g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})$ contains $\mbox{\rm Im}\ (g_{\ell})_{\mathfrak{m}_{\ell}}$ and since $(f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto, $\mbox{\rm Im}\ (g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}+V_{\mathfrak{m}_{\ell}}J(\Lambda_{\mathfrak{m}_{\ell}})=V_{\mathfrak{m}_{\ell}}$. Nakayama’s Lemma gives that $(g_{\ell}+f_{\ell})_{\mathfrak{m}_{\ell}}$ is onto and therefore there exists $T_{\ell}$ finitely generated submodule of $P$ containing $T_{\ell}^{\prime}$ such that $((g_{\ell}+f_{\ell}))_{\mathfrak{m}_{\ell}}((T_{\ell})_{\mathfrak{m}_{\ell}})=V_{\mathfrak{m}_{\ell}}$, that is $(iii)$ holds if $i=\ell$.
Theorem 6.16.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is a simple artinian ring. Then $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\cong V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$.
Proof.  Propositions 6.11 and 6.15 show that $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$. The symmetry of the hypothesis implies that $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$. By Remark 6.9, we can deduce from Corollary 2.12 that both monoids are isomorphic.
Hinohara [hinohara] proved that every projective module over an $h$-local (or more generally, weakly noetherian) domain is either finitely generated or free. We can extend this result to our setting.
Corollary 6.17.
[hinohara] Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. For each maximal ideal $\mathfrak{m}$ of $R$, assume that $\Lambda_{\mathfrak{m}}$ is Morita equivalent to a local ring. Then any infinitely generated projective $\Lambda$-module that is not finitely generated is free.
Proof.  By Remark 6.9, ${\mathcal{T}}(\Lambda)=\{0,\Lambda\}$, and hence $V(\Lambda)\sqcup B(\Lambda)=V(\Lambda)\sqcup\{\Lambda^{(\omega)}\}$. The result then follows from Theorem 6.16.
7. Rank and Genus of countably generated projective modules
We continue with the same setting as in §6. Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a locally semiperfect torsion-free algebra over $R$ such that $\Lambda_{Q}$ is simple artinian. That is $\Lambda_{Q}\cong M_{k}(D)$, for $D$ a division ring.
There are two natural monoid morphisms. The first one is the one associated with tensoring with $Q$:
${\Psi\colon V^{*}(\Lambda)}$${V^{*}(\Lambda_{Q})}$
defined by $\Psi([P])=[P_{Q}]=[S^{r}]$, where $S$ is a simple $\Lambda_{Q}$-module. Then we have a monoid morphism, that we also call $\Psi$
${\Psi\colon V^{*}(\Lambda)}$${{\mathbb{N}}_{0}^{*}}$
defined by $\Psi([P])=r$ if $P$ is finitely generated and $P_{Q}\cong S^{r}$ and $\Psi([P])=\infty$ if $P$ is not finitely generated. We call $\Psi([P])$ the rank of $P$.
Associated with the localization at the different maximal ideals, we have the monoid morphism
${\Phi\colon V^{*}(\Lambda)}$${\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}V^{*}(\Lambda_{\mathfrak{m}})}$
defined by $\Phi([P])=([P_{\mathfrak{m}}])_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}$. Recall that two countably generated projective right $\Lambda$-modules $P$ and $P^{\prime}$ have the same genus if and only if $\Phi([P])=\Phi([P^{\prime}])$. We shall refer to $\Phi$ as the genus morphism.
In this section, we want to describe the image of $\Psi$ and of $\Phi$.
At this point, we have plenty of information about $V^{*}(\Lambda)$. By Theorem 6.16, $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$, by Lemma 6.6 any ideal of $\mathcal{T}(\Lambda)$ is a trace ideal of a finitely generated projective module and, by Corollary 6.10, $\mathcal{T}(\Lambda)$ is finite.
In general, there may be plenty of finitely generated projective right modules over $\Lambda$ with the same genus that are not isomorphic. For example, let $R=\Lambda$ be a Dedekind domain. Then any localization of a projective module at a maximal ideal is free. So two projective modules have the same genus if and only if they have the same rank. Certainly, two finitely generated projective modules over a Dedekind domain with the same rank are not, in general, isomorphic.
The situation changes for projective $\Lambda$-modules whose isomorphism class is in $B(\Lambda)$.
Proposition 7.1.
Let $\Lambda$ be a locally semiperfect, torsion-free algebra over an $h$-local domain $R$ such that $\Lambda_{Q}$ is simple artinian. Then $\Phi(B(\Lambda))\subseteq\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}B(\Lambda_{\mathfrak{m}})$. Moreover, if $[P]$ and $[P^{\prime}]$ are two elements in $B(\Lambda)$, then $\Phi[P]=\Phi[P^{\prime}]$ if and only if $[P]=[P^{\prime}]$.
Proof.  For this proof, we are going to use the notation of Corollary 2.7 that is, an element $[P]\in B(\Lambda)$ can be thought of as the isomorphism class of a pair $(P/PI,I)$ where $I\in\mathcal{T}(\Lambda)$ and $P/PI$ is a finitely generated projective $\Lambda/I$-module. By Proposition 6.11, for each maximal ideal $\mathfrak{m}$ of $R$, $[P_{\mathfrak{m}}]\in B(\Lambda_{\mathfrak{m}})$ and $I_{\mathfrak{m}}$ is the ideal minimal with respect to the property that $P_{\mathfrak{m}}/P_{\mathfrak{m}}J$ is finitely generated.
Hence, $\Phi([(P/PI,I)])=([(P_{\mathfrak{m}}/P_{\mathfrak{m}}I_{\mathfrak{m}},I_{\mathfrak{m}})])_{\mathfrak{m}\in\mathrm{maxSpec}\,(R)}.$ By Lemma 1.14, it follows that $\Phi$ is injective restricted to $B(\Lambda)$.
The following straightened version of the Package Deal Theorem to our setting will be the key ingredient to describe $\Phi(V^{*}(\Lambda))=\Phi(V(\Lambda)\sqcup B(\Lambda)$.
Proposition 7.2.
Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $S$ be a simple right $\Lambda_{Q}$-module.
Fix $r\geq 1$. For each maximal ideal $\mathfrak{m}$ of $R$, let $P(\mathfrak{m})$ be a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module such that $P(\mathfrak{m})_{Q}\cong S^{r}$. Then there exists a finitely generated projective right $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong P(\mathfrak{m})$ for every maximal ideal $\mathfrak{m}$ of $R$.
Notice that, by definition, $\Psi([P])=r$.
Proof.  Let $k$ be such that $\Lambda_{Q}\cong S^{k}$. By Lemma 6.8, there is a finite set of maximal ideals of $R$, $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$, such that $\Lambda_{\mathfrak{m}}\cong M_{k}(T_{\mathfrak{m}})$ for some local ring $T_{\mathfrak{m}}$ whenever $\mathfrak{m}$ is a maximal ideal not in $A=\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$. Note that if $\mathfrak{m}\not\in A$ then $\Lambda_{\mathfrak{m}}$ is isomorphic to a ring of $k\times k$ matrices over a local ring, so it has a unique indecomposable projective module $X(\mathfrak{m})$ such that $X(\mathfrak{m})_{Q}\cong S$ and any projective $\Lambda(\mathfrak{m})$-module is isomorphic to a suitable number of copies of $X(\mathfrak{m})$. Therefore, for any $\mathfrak{m}\not\in A$, $P(\mathfrak{m})\cong X(\mathfrak{m})^{r}$.
Let $q$ be a multiple of $k$ bigger than $r$ and let $F=\Lambda^{q/k}$. Further let $U$ be a finitely generated submodule of $F$ such that $U_{Q}\cong S^{r}$. Since the embedding $U\hookrightarrow F$ splits over $\Lambda_{Q}$, it splits over almost all maximal ideals of $R$ [AHP, Lemma 4.2]. Then $U$ is a finitely generated $\Lambda$-module which is torsion-free as an $R$-module, such that $U_{\mathfrak{m}}$ is projective for almost all maximal ideals $\mathfrak{m}$ of $R$ and $U_{Q}\cong S^{r}$. So let $B$ be a finite set of maximal ideals of $R$ such that $A\subseteq B$, and satisfying that $U_{\mathfrak{m}}$ is projective for any $\mathfrak{m}\not\in B$. In particular, for any $\mathfrak{m}\not\in B$, $U_{\mathfrak{m}}\cong P(\mathfrak{m})\cong X(\mathfrak{m})^{r}$.
By Lemma 5.6, for any $\mathfrak{m}\in B$ we can choose $U(\mathfrak{m})$ a submodule of $U_{\mathfrak{m}}$ such that $U(\mathfrak{m})\cong P(\mathfrak{m})$. For $\mathfrak{m}\not\in B$ we set $U(\mathfrak{m})=U_{\mathfrak{m}}$.
Note that for any maximal ideal $\mathfrak{m}$ of $R$, $U(\mathfrak{m})_{Q}=U_{Q}$ since all these modules have length $r$ over $\Lambda_{Q}$. Therefore, we can apply Theorem 5.9 and conclude that there exists a finitely generated $\Lambda$-module $P$ of $U$ such that $P(\mathfrak{m})\cong U(\mathfrak{m})=P_{\mathfrak{m}}$ for every $\mathfrak{m}\in\operatorname{mSpec}R$. By Corollary 5.4, $P$ is a projective module over $\Lambda$.
Now we introduce the notation needed to explain in detail how the image of the genus map $\Phi$ and of the rank map $\Psi$ is.
Notation 7.3.
Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Let $S$ be a simple $\Lambda_{Q}$-module, and let $k$ be the length of $\Lambda_{Q}$, i.e., $\Lambda_{Q}\cong S^{k}$ or, equivalently, $\Psi(\Lambda)=k$. By Lemma 6.8, there is a finite set of maximal ideals $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ such that $\Lambda_{\mathfrak{m}}\cong M_{k}(T_{\mathfrak{m}})$ for some local ring $T_{\mathfrak{m}}$ whenever $\mathfrak{m}$ is a maximal ideal not in $\mathcal{S}=\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$.
If $\ell=0$, then the assumptions of Corollary 6.17 are satisfied. So in this case, the image of the rank map $\Psi$ is ${\mathbb{N}}_{0}^{*}$ and the genus map composed with the natural isomorphism $V^{*}(\Lambda_{\mathfrak{m}})\cong{\mathbb{N}}_{0}^{*}$
${\Phi\colon V^{*}(\Lambda)}$${\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}V^{*}(\Lambda_{\mathfrak{m}})\cong\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}{\mathbb{N}}_{0}^{*}}$
just sends $[P]$ to the constant sequence with value the rank of $P$.
So from now on we will assume that $\ell\geq 1$.
For $i=1,\dots,\ell$ let $U_{i,1},\dots,U_{i,t_{i}}$ be representatives of indecomposable projective right $\Lambda_{\mathfrak{m}_{i}}$-modules, and let $r_{i,1},\dots,r_{i,t_{i}}$ be the lengths of their localizations over $\Lambda_{Q}$, i.e., $(U_{i,j})_{Q}\cong S^{r_{i,j}}$.
Note that for any $i=1,\dots,\ell$, if $\Lambda_{\mathfrak{m}_{i}}$ is isomorphic, as a right $\Lambda_{\mathfrak{m}_{i}}$-module, to $U_{i,1}^{m_{i,1}}\oplus\dots\oplus U_{i,t_{i}}^{m_{i,t_{i}}}$, then we should have the relation
$\sum_{j=1}^{t_{i}}r_{i,j}m_{i,j}=(r_{i,1},\dots,r_{i,t_{i}})\begin{pmatrix}m_{i,1}\\ \vdots\\ m_{i,t_{i}}\end{pmatrix}=k.$
Now we construct examples of locally semiperfect algebras over an $h$-local domain with an infinite number of maximal ideals, showing there are no other restrictions on the parameters $k,\ell,t_{i},r_{ij},m_{ij}$ than the $\ell$ equations above.
Examples 7.4.
(1) Let $k$ be a positive integer and let $p$ be a prime. Assume we are given positive integers $r_{1},\dots,r_{t}$, $m_{1},\dots,m_{t}$ such that $\sum_{i=1}^{t}m_{i}r_{i}=k$. Our aim is to construct a semiperfect $\mathbb{Z}_{(p)}$-order $\Lambda$ in ${\rm M}_{k}(\mathbb{Q})$ such that $\Lambda\cong\bigoplus_{i=1}^{t}U_{i}^{m_{i}}$ as right modules, where $U_{1},\dots,U_{t}$ is a representative set of indecomposable projective $\Lambda$-modules and $r_{i}$ is the length of $(U_{i})_{Q}$.
Write $\{1,\dots,k\}$ as a disjoint union of sets $\bigcup_{i=1}^{t}(A_{i,1}\cup A_{i,2}\cup\cdots\cup A_{i,m_{i}})$, where $|A_{i,j}|=r_{i}$. For any $i$ and any $1\leq a,b\leq m_{i}$ we fix a bijection $f_{i,a,b}\colon A_{i,a}\to A_{i,b}$ such that $f_{i,a,a}$ is the identity of $A_{i,a}$ and $f_{i,b,c}f_{i,a,b}=f_{i,a,c}$ if $1\leq a,b,c\leq m_{i}$.
For elements $i,a,b$, $1\leq i\leq t$, $1\leq a,b\leq m_{i}$ define $e_{i,a,b}\in{\rm M}_{k}(\mathbb{Q})$ by $e_{i,a,b}=\sum_{j\in A_{i,a}}E_{f_{i,a,b}(j),j}$, where $E_{x,y}$ is a matrix unit in ${\rm M}_{k}(\mathbb{Q})$. Note that $e_{i,a,a}^{2}=e_{i,a,a}$ and $e_{i,b,c}e_{i,a,b}=e_{i,a,c}$. Further note that $e_{i,a,b}e_{j,c,d}=0$ if $1\leq i\neq j\leq k$, $1\leq a,b\leq m_{i}$ and $1\leq c,d\leq m_{j}$ or if $i=j$ and $1\leq d\neq a\leq m_{i}$.
Consider $J={\rm M}_{k}(p\mathbb{Z}_{(p)})$ and $\Lambda=J+\sum_{i=1}^{t}\sum_{1\leq a,b\leq m_{i}}e_{i,a,b}\mathbb{Z}_{(p)}$. The following facts are easily verified:
  1. (i)
    $\Lambda$ is a subring of ${\rm M}_{k}(\mathbb{Q})$ and $\Lambda_{Q}={\rm M}_{k}(\mathbb{Q})$.
  2. (ii)
    $J$ is an ideal of $\Lambda$, $J\subseteq J(\Lambda)$.
  3. (iii)
    For any $i,j$ $(e_{i,j,j}\Lambda)/(e_{i,j,j}J)$ is a simple $\Lambda$-module.
  4. (iv)
    $e_{i,a,a}\Lambda\cong e_{j,b,b}\Lambda$ if and only if $i=j$
  5. (v)
    $e_{i,a,a}\Lambda e_{j,b,b}=e_{i,a,a}Je_{j,b,b}$ whenever $1\leq i\neq j\leq t$, $1\leq a\leq m_{i},1\leq b\leq m_{j}$
From $(iii)$ it follows that $\Lambda/J$ is a direct sum of simple modules. Hence $\Lambda/J$ is semisimple artinian and therefore $J=J(\Lambda)$. Also note that the canonical projection $e_{i,j,j}\Lambda\to e_{i,j,j}\Lambda/e_{i,j,j}J$ is a projective cover of $e_{i,j,j}\Lambda/e_{i,j,j}J$. So $\Lambda$ is semiperfect and $e_{i,j,j}\Lambda$ are indecomposable projectives. Also $e_{i,a,a}\Lambda\cong e_{j,b,b}\Lambda$ if and only if $i=j$. So we have that $\Lambda=\bigoplus_{i=1}^{t}\bigoplus_{a=1}^{m_{i}}e_{i,a,a}\Lambda\cong\bigoplus_{i=1}^{t}(e_{i,1,1}\Lambda)^{m_{i}}$. Also it is easy to check that $e_{i,a,a}{\rm M}_{k}(\mathbb{Q})$ has length $r_{i}$ since $|A_{i,a}|=r_{i}$.
(2) Let $k,\ell$ be positive integers $p_{1},\dots,p_{\ell}$ pairwise different primes. Assume that for each $i$ positive integers $t_{i}$, $r_{i,j},m_{i,j}$, $j\in\{1,\dots,t_{i}\}$ are given such that $\sum_{j=1}^{t_{i}}r_{i,j}m_{i,j}=k$ for every $i=1,\dots,\ell$. By part (a) there exists a semiperfect $\mathbb{Z}_{(p_{i})}$-order $\Lambda(p_{i})$ in ${\rm M}_{k}(\mathbb{Q})$ such that $\Lambda(p_{i})\cong\bigoplus_{j=1}^{t_{i}}U_{i,j}^{m_{i,j}}$ where $U_{i,1},\dots U_{i,t_{i}}$ are pairwise non-isomorphic projective $\Lambda(p_{i})$-modules and $(U_{i,j})_{Q}$ has rank $r_{i,j}$.
Now the package deal theorem finds a $\mathbb{Z}$-submodule $\Lambda\subseteq{\rm M}_{k}(\mathbb{Z})$ such that $\Lambda_{(p_{i})}=\Lambda(p_{i})$ for $i=1,\dots,\ell$ and $\Lambda_{(q)}={\rm M}_{k}(\mathbb{Z}_{(q)})$ if $q$ is a prime different from $p_{1},\dots,p_{\ell}$. It is easy to see that $\Lambda$ has to be a subring of ${\rm M}_{k}(\mathbb{Z})$ and that $\Lambda_{(p_{i})}\simeq\Lambda(p_{i})$ as rings.
Remark 7.5.
We keep the Notation 7.3. First we observe that
${\Phi\colon V^{*}(\Lambda)}$${\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}V^{*}(\Lambda_{\mathfrak{m}})=\prod_{i=1}^{\ell}V^{*}(\Lambda_{\mathfrak{m}_{i}})\times\prod_{\mathfrak{m}\not\in\mathcal{S}}V^{*}(\Lambda_{\mathfrak{m}})}$
So for any $\mathfrak{m}\not\in\mathcal{S}$, $V^{*}(\Lambda_{\mathfrak{m}})\cong{\mathbb{N}}_{0}^{*}$ and, for any $[P]\in V^{*}(\Lambda)$, this isomorphism sends $[P_{\mathfrak{m}}]$ to its rank which is $\Psi[P]$. So the genus map is completely determined by its image on $V^{*}(\Lambda_{\mathfrak{m}})=\prod_{i=1}^{\ell}V^{*}(\Lambda_{\mathfrak{m}_{i}})$.
Then there is a monoid morphism
${\Phi^{\prime}\colon V(\Lambda)}$${\prod_{i=1}^{\ell}V(\Lambda_{\mathfrak{m}_{i}})\cong\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)}$
$\Phi^{\prime}([P])=([P_{\mathfrak{m}_{1}}],\dots,[P_{\mathfrak{m}_{\ell}}])=\Big{(}\textstyle\sum_{j=1}^{t_{1}}[U_{1,j}]^{x_{1,j}},\dots,\sum_{j=1}^{t_{\ell}}[U_{\ell,j}]^{x_{\ell,j}}\Big{)}.$
Where the last isomorphism assigns to $\Phi^{\prime}([P])$ the element
$((x_{1,1},\dots,x_{1,t_{1}}),\dots,(x_{\ell,1},\dots,x_{\ell,t_{\ell}}))\in\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right).$
If $P_{Q}\cong S^{r}$, then for any $i=1,\dots,\ell$, we have the identity
$\sum_{j=1}^{t_{i}}r_{i,j}x_{i,j}=(r_{i,1},\dots,r_{i,t_{i}})\begin{pmatrix}x_{i,1}\\ \vdots\\ x_{i,t_{i}}\end{pmatrix}=r$ (10)
Conversely, by Proposition 7.2, an element $\mathbf{X}\in\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$ corresponds to a finitely generated projective right $\Lambda$-module if and only if there exists $r$ such that the components of $\mathbf{X}$ satisfy the system of equations (10). So this describes all the possible genus of finitely generated projective $\Lambda$-modules.
So we have proved the following.
Proposition 7.6.
Let $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$ in $\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$. Then $\mathbf{X}\in\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$ is an element of $\mathcal{A}$ if and only if there exists $r\in{\mathbb{N}}_{0}$ such that the components of $\mathbf{X}$ satisfy
$\sum_{j=1}^{t_{i}}r_{i,j}x_{i,j}=(r_{i,1},\dots,r_{i,t_{i}})\begin{pmatrix}x_{i,1}\\ \vdots\\ x_{i,t_{i}}\end{pmatrix}=r$
for any $i=1,\dots,\ell$.
If $P$ is a finitely generated projective right $\Lambda$-module such that $\Phi^{\prime}([P])=\mathbf{X}$, then $P_{Q}\cong S^{r}$. So
[…]
Remark 7.7.
If $\ell=1$, and $U_{1},\dots,U_{t}$ are the indecomposable projective modules of $\Lambda_{\mathfrak{m}}$, up to isomorphism, and $\Psi(U_{i})=r_{i}$ for $i=1,\dots,t$. Then $\Psi(V(\Lambda))$ is the submonoid of ${\mathbb{N}}_{0}$ generated by $r_{1},\dots,r_{\ell}$. By Proposition 7.2, $\Phi^{\prime}(V(R))\cong{\mathbb{N}}_{0}^{t}$, because, for any $i=1,\dots,\ell$, there exists an (indecomposable) finitely generated projective right $\Lambda$-module $P_{i}$ such that $(P_{i})_{\mathfrak{m}_{i}}\cong U_{i}$.
Clearly, $\Psi(V^{*}(\Lambda))=\Psi(V(\Lambda))\cup\{\infty\}$ and $\Phi^{\prime}(V^{*}(\Lambda))\cong({\mathbb{N}}_{0}^{*})^{t}$. Notice that, in this case, by Proposition 7.1, any projective $\Lambda$-module is isomorphic to a direct sum of finitely generated projective $\Lambda$-modules.
Now we proceed to extend this description of the genus to countably generated projective modules. In view of Remark 7.7, we may assume $\ell\geq 2$.
Proposition 7.8.
Extend the monoid morphism $\Phi^{\prime}$ to $V^{*}(\Lambda)$ and enlarge ${\mathbb{N}}_{0}$ to ${\mathbb{N}}_{0}^{*}={\mathbb{N}}_{0}\cup\{\infty\}$. Then $\mathbf{Y}\in\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$ is an element in $\Phi^{\prime}(B(\Lambda))=\Phi^{\prime}(V^{*}(\Lambda)\setminus V(\Lambda))$ if and only if its components are solutions of
$\sum_{j=1}^{t_{i}}r_{i,j}y_{i,j}=(r_{i,1},\dots,r_{i,t_{i}})\begin{pmatrix}y_{i,1}\\ \vdots\\ y_{i,t_{i}}\end{pmatrix}=\infty$ (11)
for any $i=1,\dots,\ell$.
Proof.  By Theorem 6.16, $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$, so $\Phi^{\prime}(B(\Lambda))=\Phi^{\prime}(V^{*}(\Lambda)\setminus V(\Lambda))$.
Let $\mathbf{Y}\in\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$. For $i=1,\dots,\ell$ let $S_{i}(\mathbf{Y})=\{j\mid y_{i,j}=\infty\}$. Observe that $\mathbf{Y}$ is a solution of the equations (11) if and only if $S_{i}(\mathbf{Y})\neq\emptyset$ for any $i=1,\dots,\ell$.
Let $P$ be a countably generated right $\Lambda$-module such that it is not finitely generated. Then $\Psi(P)=\infty$, which implies that, for any $i=1,\dots,\ell$, there exists $j$ such that the indecomposable module $U_{i,j}$ appears infinitely many times in the decomposition of $P_{\mathfrak{m}}$ as a direct sum of indecomposable modules. Therefore, if $\mathbf{Y}=\Phi^{\prime}([P])$ then $S_{i}(\mathbf{Y})\neq\emptyset$ for any $i=1,\dots,\ell$. So $\mathbf{Y}$ is a solution of the system (11).
To prove the converse, let $\mathbf{Y}\in\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$ such that $S_{i}(\mathbf{Y})\neq\emptyset$ for any $i=1,\dots,\ell$. Then we show that there is a countably generated projective module $P$ such that $\mathbf{Y}=\Phi^{\prime}([P])$.
For $i=1,\dots,\ell$, let $P^{\prime}(\mathfrak{m}_{i})=\bigoplus_{j\in S_{i}(\mathbf{Y})}U_{i,j}$ and let $r_{i}$ be the rank of $P^{\prime}(\mathfrak{m}_{i})$. Let $r$ be the least common multiple of $r_{1},\dots,r_{\ell},k$. For $i=1,\dots,\ell$, set $P(\mathfrak{m}_{i})=P^{\prime}(\mathfrak{m}_{i})^{\frac{m}{r_{i}}}$ and set $P(\mathfrak{m})=\Lambda_{\mathfrak{m}}^{\frac{r}{k}}$. By Package deal Theorem 5.12, there exists a finitely generated projective $\Lambda$-module $P^{\prime}$ such that $P^{\prime}_{\mathfrak{m}}\cong P(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$.
Let $I$ be the trace ideal of $P^{\prime}$. Then $I_{\mathfrak{m}_{i}}=\sum_{j\in S_{i}(\mathbf{Y})}\mathrm{Tr}(U_{i,j})$ for $i=1,\dots,\ell$, and $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ for the other maximal ideals of $R$. As $R$ is $h$-local,
$\Lambda/I\cong\prod_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}.$
For any $i=1,\dots,\ell$, set $A_{i}=\{1,\dots,t_{i}\}\setminus S_{i}(\mathbf{Y})$. Then the indecomposable projective modules over $\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$ are $\{U_{i,j}/U_{i,j}I_{\mathfrak{m}_{i}}\}_{j\in A_{i}}$.
Now consider the finitely generated projective $\Lambda/I$-module
$\overline{P}=\bigoplus_{i=1}^{\ell}\Big{(}\bigoplus_{j\in A_{i}}(U_{i,j}/U_{i,j}I_{\mathfrak{m}_{i}})^{y_{i,j}}\Big{)}.$
By Lemma 1.9, there exists a countably generated projective $\Lambda$-module $P$ such that $P/PI\cong\overline{P}$. Then $P\oplus(P^{\prime})^{(\omega)}$ is an $I$-big countably generated projective $\Lambda$-module such that $\Psi^{\prime}(P\oplus(P^{\prime})^{(\omega)})=\mathbf{Y}.$
8. Infinite sums of finitely generated projective modules
In this section we keep working with an algebra $\Lambda$ over an $h$-local domain $R$, with field of fractions $Q$ and such that $\Lambda_{Q}$ is simple artinian.
We keep using the notation of the previous section. In particular, the one introduced in Notation 7.3. As in Proposition 7.6, let $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$. If $\ell$ is either 00 or $1$, then every projective module is a direct sum of finitely generated ones (cf. Remark 7.7). So we assume $\ell\geq 2$.
Let $\mathcal{B}$ be the image of $B(\Lambda)=V^{*}(\Lambda)\setminus V(\Lambda)$ in $\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$.
By Proposition 7.1, the genus classifies the modules of $B(\Lambda)$ up to isomorphism. Therefore, the modules in $B(\Lambda)$ that are isomorphic to a direct sum of finitely generated modules are the ones such that the image of its isomorphism class via $\Phi^{\prime}$ lies in
$\mathcal{A}+\infty\cdot(\mathcal{A}\setminus\{0\})=\{a+\infty\cdot a^{\prime}\mid a,a^{\prime}\in\mathcal{A}\mbox{ and }a^{\prime}\neq 0\}.$
In this section we will examine the question whether any projective $\Lambda$-module is a direct sum of finitely generated ones. By Proposition 2.8, we only need to characterize when finitely generated projective modules modulo an ideal of $\mathcal{T}(\Lambda)$ lift to finitely generated projective modules. Our next Lemma shows that we only need to ensure this property for the minimal elements in $\mathcal{T}(\Lambda)\setminus\{0\}$.
Lemma 8.1.
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. The following are equivalent
  1. (i)
    For any $I\in\mathcal{T}(\Lambda)$ and any finitely generated projective $\Lambda/I$-module $P^{\prime}$ there exists a finitely generated projective module $\Lambda$-module $P$ such that $P^{\prime}\cong P/PI$.
  2. (ii)
    For any $I$ which is a minimal element of $\mathcal{T}(\Lambda)\setminus\{0\}$ and any finitely generated projective $\Lambda/I$-module $P^{\prime}$ there exists a finitely generated projective module $\Lambda$-module $P$ such that $P^{\prime}\cong P/PI$.
Proof.  Of course, only $(ii)\Rightarrow(i)$ needs a proof. To prove $(ii)\Rightarrow(i)$, let $I$ be a nonzero trace of $\Lambda$ containing a minimal trace $X\in T(\Lambda)\setminus\{0\}$. Such an $X$ exists by Corollary 6.10.
As $X_{Q}=\Lambda_{Q}$, $\Lambda/X$ is a torsion $R$-module. As $R$ is $h$-local, the canonical homomorphism $\Lambda/X\to\prod_{\mathfrak{m}\in\operatorname{mSpec}R}(\Lambda/X)_{\mathfrak{m}}$ induces an isomorphism $\varphi\colon\Lambda/X\to\bigoplus_{\mathfrak{m}\in\operatorname{mSpec}R}(\Lambda/X)_{\mathfrak{m}}$. Notice that, by Lemma 5.8, $X_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$.
Since $(\Lambda/X)_{\mathfrak{m}}\cong\Lambda_{\mathfrak{m}}/X_{\mathfrak{m}}$ is a semiperfect ring, the module $\bigoplus_{\mathfrak{m}\in\operatorname{mSpec}R}(\Lambda/X)_{\mathfrak{m}}$ can be seen as a finite product of semiperfect rings and $\varphi$ is an isomorphism of rings. In particular, $\Lambda/X$ is semiperfect.
Let $P^{\prime}$ be a finitely generated projective $\Lambda/I$-module. By Lemma 6.1, there exists a finitely generated projective $\Lambda/X$-module $P_{1}$ such that $P_{1}/P_{1}I\cong P^{\prime}$. Now we can use $(ii)$ to find a finitely generated projective $\Lambda$-module $P$ such that $P/PX=P_{1}$. Then $P/PI\cong(P/PX)/(PI/PX)\cong(P_{1}/P_{1}I)\cong P^{\prime}$.
Next Lemma is going to be useful to be able to check when the criteria of Lemma 8.1 is satisfied.
Lemma 8.2.
Let $2\leq\ell\in\mathbb{N}$. Let $r_{1},\dots,r_{\ell},s\in\mathbb{N}$. Let $m=\mathrm{lcm}(r_{1},\dots,r_{\ell-1})>0$. Then the system
$\begin{pmatrix}r_{1}&0&\dots&0&-s\\ 0&r_{2}&\dots&0&-s\\ \vdots&&\ddots&&\vdots\\ 0&0&\dots&r_{\ell-1}&-s\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{\ell}\end{pmatrix}=\begin{pmatrix}r_{\ell}\\ \vdots\\ r_{\ell}\end{pmatrix}$ (12)
has a solution in $\mathbb{N}$ if and only if the equation $my-sx_{\ell}=r_{\ell}$ has a solution in $\mathbb{N}$. If $(y,s)$ is such a solution, then
$\left(x_{1}=\frac{m}{r_{1}}y,\dots,x_{\ell-1}=\frac{m}{r_{\ell-1}}y,x_{\ell}\right)$
is a solution of (12). Therefore, (12) has a solution in $\mathbb{N}$ if and only if $\gcd(m,s)$ divides $r_{\ell}$.
Proof.  The claim follows because
$\begin{pmatrix}1&0&\dots&0\\ -1&1&\dots&0\\ \vdots&&\ddots&\vdots\\ -1&0&\dots&1\end{pmatrix}\begin{pmatrix}r_{1}&0&\dots&0&-s\\ 0&r_{2}&\dots&0&-s\\ \vdots&&\ddots&&\vdots\\ 0&0&\dots&r_{\ell-1}&-s\end{pmatrix}=\begin{pmatrix}r_{1}&0&\dots&0&-s\\ -r_{1}&r_{2}&\dots&0&0\\ \vdots&&\ddots&&\vdots\\ -r_{1}&0&\dots&r_{\ell-1}&0\end{pmatrix}$
so the system (12) is equivalent to the system
$\begin{pmatrix}r_{1}&0&\dots&0&-s\\ -r_{1}&r_{2}&\dots&0&0\\ \vdots&&\ddots&&\vdots\\ -r_{1}&0&\dots&r_{\ell-1}&0\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{\ell}\end{pmatrix}=\begin{pmatrix}r_{\ell}\\ 0\\ \vdots\\ 0\end{pmatrix}$
which has a solution in $\mathbb{N}$ if and only if the equation $my-sx_{\ell}=r_{\ell}$ has a solution in $\mathbb{N}$. Also, it is clear that the solutions are related as claimed in the statement.
Proposition 8.3.
Under the Notation 7.3, the following are equivalent
  1. (i)
    Every projective $\Lambda$-module is a direct sum of finitely generated modules.
  2. (ii)
    For every choosing $a_{1},a_{2},\dots,a_{\ell}$, $1\leq a_{i}\leq t_{i}$, and for every $j\in\{1,\dots,\ell\}$ and $1\leq b\neq a_{j}\leq t_{j}$ there are non-negative integers $x_{1},x_{2},\dots,x_{\ell}$ such that $r_{j,b}+x_{j}r_{j,a_{j}}=x_{i}r_{i,a_{i}}$ for every $i\neq j$.
  3. (iii)
    For every choosing $a_{1},a_{2},\dots,a_{\ell}$, $1\leq a_{i}\leq t_{i}$, and for every $j\in\{1,\dots,\ell\}$ and $1\leq b\neq a_{j}\leq t_{j}$ then $\gcd(\mathrm{lcm}(\{r_{1,a_{1}},\dots,r_{\ell,a_{\ell}}\}\setminus\{r_{j,a_{j}}\}),r_{j,a_{j}})$ divides $r_{j,b}$.
Proof.  Assume $(i)$ is true, so finitely generated projective modules of $\Lambda$ modulo a (minimal) element of $\mathcal{T}(\Lambda)\setminus\{0\}$ can be lifted to finitely generated projective modules $\Lambda$-modules.
By Remark 6.9, $\Lambda$ has an ideal $I\in\mathcal{T}(\Lambda)$ such that $I_{\mathfrak{m}_{i}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}_{i}}}(U_{i,a_{i}})$ for each $i=1,\dots,\ell$. As $\Lambda_{Q}$ is simple artinian and $R$ is $h$-local, $\Lambda/I\cong\bigoplus_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$. Therefore, every (finitely generated) projective $\Lambda/I$-module is of the form $\bigoplus_{i=1}^{\ell}M_{i}$, where $M_{i}$ is a (finitely generated) projective $(\Lambda/I)_{\mathfrak{m}_{i}}$-module. Note that $(\Lambda/I)_{\mathfrak{m}_{i}}\cong\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$ which is a semiperfect ring, so projective modules over it are described in Proposition 6.2 (iv).
Consider a projective $\Lambda/I$-module $P^{\prime}$ such that $P_{\mathfrak{m}_{j}}^{\prime}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$ and $P_{\mathfrak{m}_{i}}^{\prime}=0$ if $i\neq j$. By assumption, there exists $P$ a finitely generated projective $\Lambda$-module such that $P/PI\cong P^{\prime}$. Then $P_{\mathfrak{m}_{i}}\cong U_{i,a_{i}}^{x_{i}}$ if $i\neq j$ and $P_{\mathfrak{m}_{j}}\cong U_{j,b}\oplus U_{j,a_{j}}^{x_{j}}$. Since $(P_{\mathfrak{m}_{i}})_{Q}\cong P_{Q}\cong(P_{\mathfrak{m}_{j}})_{Q}$, we get $x_{i}r_{i,a_{i}}=r_{j,b}+x_{j}r_{j,a_{j}}$. So $(ii)$ holds.
Conversely, let $I$ be a minimal nonzero trace of $\Lambda$. Then $I_{\mathfrak{m}_{i}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}_{i}}}(U_{i,a_{i}})$ for some $a_{i}$, $i\in\{1,\dots,\ell_{i}\}$. Recall that $\Lambda/I\cong\bigoplus_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$ is a semiperfect ring, hence it is sufficient to lift an indecomposable projective $\Lambda/I$-module to a finitely generated projective $\Lambda$-module. So let $P^{\prime}$ be a finitely generated indecomposable projective $\Lambda/I$-module. Regarding the decomposition of $\Lambda/I$, $P^{\prime}_{\mathfrak{m}_{i}}=0$ for all $i$’s with exactly one exception. Assume $P_{\mathfrak{m}_{j}}^{\prime}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$, where $b\neq a_{j}$. Take $x_{1},\dots,x_{\ell}\in\mathbb{N}_{0}$ such that the equations in $(ii)$ are satisfied. Let $r=r_{j,b}+x_{j}r_{j,a_{j}}$.
For $i=1,\dots,\ell$ set $U(\mathfrak{m}_{i})=U_{i,a_{i}}^{x_{i}}$ if $i\neq j$ and $U(\mathfrak{m}_{j})=U_{j,b}\oplus U_{j,a_{j}}^{x_{j}}$. For $\mathfrak{n}\in\operatorname{mSpec}(R)\setminus\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$ let $U(\mathfrak{n})$ be a projective $\Lambda_{\mathfrak{n}}$-module such that $U(\mathfrak{n})_{Q}$ has length $r$ (it exists by Lemma 6.8). Because of the choice of $x_{1},\dots,x_{\ell}$, we can use Proposition 7.2 to see the existence of a finitely generated projective $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong U(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. Then $P_{\mathfrak{m}_{i}}I_{\mathfrak{m}_{i}}=P_{\mathfrak{m}_{i}}$ if $i\neq j$ and $P_{\mathfrak{m}_{j}}/P_{\mathfrak{m}_{j}}I_{\mathfrak{m}_{j}}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$. Therefore $P/PI\cong\bigoplus_{i=1}^{\ell}P_{\mathfrak{m}_{i}}/P_{\mathfrak{m}_{i}}I_{\mathfrak{m}_{i}}\cong P^{\prime}$. This proves $(ii)\Rightarrow(i)$.
The equivalence of $(ii)$ and $(iii)$ follows from Lemma 8.2.
Example 8.4.
Under the assumptions in Notation 7.3, let $\ell=2$, and assume that $U_{1,1},U_{1,2}$ and $U_{2,1},U_{2,2}$ are representatives of the finitely generated indecomposable projective right modules over $\Lambda_{\mathfrak{m}_{1}}$ and $\Lambda_{\mathfrak{m}_{2}}$, respectively. Suppose that
$r_{1,1}=2,\quad r_{1,2}=4,\quad r_{2,1}=3,\quad r_{2,2}=9.$
Notice that a locally semiperfect algebra $\Lambda$ exists by Example 7.4.
Since
$12=r_{1,1}2+r_{1,2}4=r_{2,1}+r_{2,1},$
$\Lambda$ can be taken to be an algebra over an $h$-local domain $R$ with infinitely many maximal ideals, such that
  • (i)
    $\Lambda_{Q}\cong M_{12}(T_{Q})$;
  • (ii)
    $\Lambda_{\mathfrak{m}_{1}}$ is a semiperfect ring such that modulo its Jacobson radical is isomorphic to $M_{2}(T_{1,1})\times M_{2}(T_{1,2})$, and $\Lambda_{\mathfrak{m}_{2}}$ is a semiperfect ring that is a product of two division rings modulo the Jacobson radical;
  • (iii)
    for any other maximal ideal $\mathfrak{m}$ of $R$ the localization $\Lambda_{\mathfrak{m}}\cong M_{12}(T_{\mathfrak{m}})$ where $T_{\mathfrak{m}}$ is a local ring.
By Proposition 7.6, to compute $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$ in $\mathbb{N}_{0}^{2}\times\mathbb{N}_{0}^{2}$ we need to find the $((x_{1,1},x_{1,2}),(x_{2,1},x_{2,2}))\in\mathbb{N}_{0}^{2}\times\mathbb{N}_{0}^{2}$ which are solutions of the system of equations:
$\begin{split}2x_{1,1}+4x_{1,2}&=r\\ 3x_{2,1}+9x_{2,2}&=r,\end{split}$ (13)
for some fixed $r\in\mathbb{N}_{0}$.
Since $\mathrm{lcm}(\gcd(2,4),\gcd(3,9))=6\mid r$, the image of the rank function $\Psi$ lies in $6\mathbb{N}_{0}$ and, since there are finitely generated projective modules of rank $6$, in fact $\Psi(V(\Lambda))=6\mathbb{N}_{0}$.
As each $r_{i,j}$ is coprime to $r_{i^{\prime},j^{\prime}}$ for every $i^{\prime}\neq i$ and $j^{\prime}\neq j$, statement (iii) in Proposition 8.3 is satisfied. So every projective $\Lambda$-module is a direct sum of finitely generated modules.
Denote an element in $\mathcal{A}$ by $((x_{1,1},x_{1,2}),(x_{2,1},x_{2,2}))_{r}$, where $x_{1,1},x_{1,2},x_{2,1},x_{2,2},r\in\mathbb{N}_{0}$ satisfy (13). Hence, a minimal set of generators of $\mathcal{A}$ is
$((3,0),(2,0))_{6},\quad((1,1),(2,0))_{6},\quad((0,3),(4,0))_{12},$
$((6,0),(1,1))_{12},\quad((4,1),(1,1))_{12},\quad((2,2),(1,1))_{12},\quad((0,3),(1,1))_{12},$
$((9,0),(0,2))_{18},\quad((7,1),(0,2))_{18},\quad((5,2),(0,2))_{18},$
$((3,3),(0,2))_{18},\quad((1,4),(0,2))_{18},\quad((0,9),(0,4))_{36}.$
So the isomorphism class of finitely generated projective $\Lambda$-module has as an image in $\mathcal{A}$ a non-negative integral linear combination of elements in this generating set.
If we extend the monoid morphism $\Phi^{\prime}$ to $V^{*}(\Lambda)$ and enlarge $\mathbb{N}_{0}$ to $\mathbb{N}_{0}^{*}=\mathbb{N}_{0}\cup\{\infty\}$, then we can also compute $\mathcal{B}=\Phi^{\prime}(B(\Lambda))$ in $(\mathbb{N}_{0}^{*})^{2}\times(\mathbb{N}_{0}^{*})^{2}$. A minimal set of generators of $\mathcal{B}$ is
$((\infty,0),(\infty,0)),\quad((\infty,0),(0,\infty)),\quad((0,\infty),(\infty,0)),\quad((0,\infty),(0,\infty)),$
$((\infty,1),(\infty,0)),\quad((\infty,1),(0,\infty)),\quad((1,\infty),(\infty,0)),\quad((1,\infty),(0,\infty)),$
$((\infty,0),(\infty,1)),\quad((\infty,0),(1,\infty)),\quad((0,\infty),(\infty,1)),\quad((0,\infty),(1,\infty)).$
and, as explained before, all of them correspond to the isomorphism class of a countably generated projective module that is a direct sum of finitely generated modules. For example,
$((\infty,1),(\infty,0))=((1,1),(2,0))_{6}+\infty\cdot((3,0),(2,0))_{6}.$
Example 8.5.
Under the assumptions in Notation 7.3, there are examples of algebras $\Lambda$ where there is a projective $\Lambda$-module which is not a direct sum of finitely generated modules. Set now
$r_{1,1}=2,\quad r_{1,2}=8,\quad r_{2,1}=2,\quad r_{2,2}=4.$
By Example 7.4, we know there exist locally semiperfect algebras $\Lambda$ with such invariants.
To determine $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$ we have to solve the following system of linear equations:
$\begin{split}2x_{1,1}+8x_{1,2}&=r\\ 2x_{2,1}+4x_{2,2}&=r.\end{split}$ (14)
Since $\gcd(\{r_{1,2},r_{2,2}\}\setminus\{r_{1,2}\},r_{1,2})=\gcd(4,8)=4\nmid 2=r_{1,1}$, Proposition 8.3 (iii) fails. Therefore, in this case, there are projective $\Lambda$-modules which are not a direct sum of finitely generated modules. We are going to compute an explicit example of such a module.
As in Example 8.4, if we extend the monoid morphism $\Phi^{\prime}$ to $V^{*}(\Lambda)$ and enlarge $\mathbb{N}_{0}$ to $\mathbb{N}_{0}^{*}=\mathbb{N}_{0}\cup\{\infty\}$, then we can compute $\mathcal{B}=\Phi^{\prime}(B(\Lambda))$ in $(\mathbb{N}_{0}^{*})^{2}\times(\mathbb{N}_{0}^{*})^{2}$. The generating set is the same as in Example 8.4, because it is independent of the $r_{i,j}$’s chosen, it only depends on how many of them there are.
Let $[P]\in B(\Lambda)$ be such that $\Phi^{\prime}([P])=((0,\infty),(1,\infty))$. Then there is no $\mathbf{X}\in\mathbb{N}_{0}^{2}\times\mathbb{N}_{0}^{2}$ of the form $\mathbf{X}=((0,x_{1,2}),(1,x_{2,2}))$ being a solution of (14) since $8x_{1,2}\equiv 0\mod 2$ and $1+4x_{2,2}\equiv 1\mod 2$. Therefore, there is no finitely generated projective $\Lambda$-module $P^{\prime}$ such that $\Phi^{\prime}([P^{\prime}])=((0,x_{1,2}),(1,x_{2,2}))$, so $P$ is a countably generated projective $\Lambda$-module which is not a direct sum of finitely generated modules.
Let $P_{1}$ be a finitely generated $\Lambda$-module such that $\Phi^{\prime}([P_{1}])=((0,1),(0,2))_{8}$, and let $P_{2}$ be such that $\Phi^{\prime}([P_{2}])=((0,1),(2,4))_{8}$. By Proposition 7.1, $P\cong P\oplus P_{1}^{(\omega)}$ and $P^{2}\cong P_{2}\oplus P_{1}^{(\omega)}$. Thus $P^{2}$ is a direct sum of finitely generated projective modules.
9. Torsion-free modules over $h$-local domains
Now we are ready to apply our results to $\mbox{\rm Add}(M)$ for $M$ a finitely generated torsion-free module $M$ over an $h$-local domain. That is, we are going to apply the previous results to $\Lambda=\operatorname{End}_{R}(M)$. We need to keep in mind the following lemma.
Lemma 9.1.
[AHP, Lemma 2.18] Let $R$ be a commutative domain with field of fractions $Q$. Let $M$ be a non-zero finitely generated torsion-free module with endomorphism ring $\Lambda$. Then
  • (i)
    $\Lambda$ is a torsion-free $R$-module containing $R$.
  • (ii)
    $\Lambda_{\mathfrak{m}}\cong\operatorname{End}_{R_{\mathfrak{m}}}(M_{\mathfrak{m}})$ is a semilocal ring for every maximal ideal $\mathfrak{m}$ of $R$.
  • (iii)
    $\Lambda_{Q}\cong M_{k}(Q)$ is a simple artinian ring, where $k$ is the rank of $M$.
  • (iv)
    For any non-zero two-sided ideal $I$ of $\Lambda$, $I\cap R\neq\{0\}$.
If, in addition, $R$ is $h$-local, then
  • (v)
    $M_{\mathfrak{m}}$ is free for almost all maximal ideals of $R$.
  • (vi)
    For any non-zero two-sided ideal $I$ of $\Lambda$, there are canonical isomorphisms
    $R/(I\cap R)\longrightarrow(R/(I\cap R))_{\mathfrak{m}_{1}}\times\dots\times(R/(I\cap R))_{\mathfrak{m}_{t}}$
    and
    $\Lambda/I\longrightarrow(\Lambda/I)_{\mathfrak{m}_{1}}\times\dots\times(\Lambda/I)_{\mathfrak{m}_{t}},$
    where $\{\mathfrak{m}_{1},\dotsc,\mathfrak{m}_{t}\}$ are the unique ideals of $R$ containing $I\cap R$.
So let $M_{R}$ be a finitely generated torsion-free module of rank $k$ over an $h$-local domain $R$. Let $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ be the list of those maximal ideals of $R$ over which $M_{\mathfrak{m}}$ is not free. Assume further that $\Lambda$ is locally semiperfect, that is, that each $M_{\mathfrak{m}}$ is a direct sum of modules with local endomorphism rings. For $i=1,\dots,\ell$, let $M_{\mathfrak{m}_{i}}\cong\bigoplus_{j=1}^{t_{i}}M_{i,j}^{m_{i,j}}$, where $M_{i,1},\dots,M_{i,t_{i}}$ are pairwise non-isomorphic indecomposable modules and $m_{i,j}\in\mathbb{N}$. Note that $\Lambda_{\mathfrak{m}_{i}}\cong\bigoplus_{j=1}^{t_{i}}U_{i,j}^{m_{i,j}}$, where $U_{i,j}=\operatorname{Hom}_{R_{\mathfrak{m}_{i}}}(M_{\mathfrak{m}_{i}},M_{i,j})$, is an indecomposable decomposition of $\Lambda_{\mathfrak{m}_{i}}$. Let $r_{i,j}$ be the length of $(U_{i,j})_{Q}$. As a $Q$-module, $(U_{i,j})_{Q}=Q^{kr_{i,j}}$. On the other hand, we also have isomorphisms of $Q$-modules $(U_{i,j})_{Q}\cong\operatorname{Hom}_{Q}(M_{Q},(M_{i,j})_{Q})\cong(M_{i,j})_{Q}^{k}=Q^{k\operatorname{rank}(M_{i,j})}$. Therefore $r_{i,j}=\operatorname{rank}(M_{i,j})$.
Lemma 9.2.
[…]
  1. (i)
    $\operatorname{End}_{R}(M)$ is a locally semiperfect ring.
  2. (ii)
    for every $\mathfrak{m}\in\operatorname{mSpec}R$, $M_{\mathfrak{m}}$ contains a direct summand isomorphic to $R_{\mathfrak{m}}$.
Proof.  The assumption $\operatorname{Tr}_{R}(M)=R$ is equivalent to $R\in\mbox{\rm add}(M)$. In particular, $\mbox{\rm Add}(M)$ contains $R^{(\omega)}\oplus M$, so $(i)$ follows from [AHP, Proposition 6.1]. Note that $\operatorname{Tr}_{R_{\mathfrak{m}}}(M_{\mathfrak{m}})=R_{\mathfrak{m}}$ for any maximal ideal $\mathfrak{m}\in\operatorname{mSpec}R$, hence $R_{\mathfrak{m}}\in\mbox{\rm add}(M_{\mathfrak{m}})$. Since $R_{\mathfrak{m}}$ is local, $R_{\mathfrak{m}}$ has to be isomorphic to a direct summand of $M_{\mathfrak{m}}$.
Now we are ready to prove our main application to direct sum decompositions of modules, which complements the results in [AHP].
Theorem 9.3.
Let $R$ be an $h$-local domain and let $M$ be a non-zero finitely generated torsion-free $R$-module with endomorphism ring $\Lambda=\operatorname{End}_{R}(M)$. Consider the following statements,
  1. (i)
    $\Lambda$ is locally semiperfect and if $X$ and $Y$ are indecomposable direct summands of $M_{\mathfrak{m}}$ and $N_{\mathfrak{n}}$ for some $\mathfrak{m}\neq\mathfrak{n}\in\operatorname{mSpec}R$, then $\operatorname{rank}(X)$ and $\operatorname{rank}(Y)$ are coprime.
  2. (ii)
    Every module in $\mbox{\rm Add}(M)$ is a direct sum of finitely generated modules.
Then $(i)$ always implies $(ii)$ and the converse is true provided $M$ is a generator and $R$ has Krull dimension $1$.
Proof.  Notice that by Lemma 9.1, $\Lambda$ is torsion-free as an $R$-module, and $\Lambda_{Q}$ is simple artinian.
Let us show that $(i)\Rightarrow(ii)$. Assume that $(i)$ is true. Note that $r_{i,a}$ and $r_{j,b}$ are coprime whenever $i\neq j$, $a\in\{1,\dots,t_{i}\}$, $b\in\{1,\dots,t_{j}\}$. We want to check that $(ii)$ of Proposition 8.3 is true.
Consider $a_{1},\dots,a_{\ell}$, $a_{i}\in\{1,\dots,t_{i}\}$ and $A=\prod_{i=1}^{\ell}r_{i,a_{i}}$. Fix $1\leq j\leq\ell$. Since $r_{1,a_{1}},\dots,r_{\ell,a_{\ell}}$ are pairwise coprime, there exists $t\in\{0,\dots,A-1\}$ such that $t\equiv 0\ ({\rm mod}\ r_{i,a_{i}})$ if $i\neq j$ and $t\equiv r_{j,b}\ ({\rm mod}\ r_{j,a_{j}})$. If $t<r_{j,b}$ we can increase $t$ by a suitable multiple of $A$, so we may assume $t\geq r_{j,b}$. Set $x_{i}:=t/r_{i,a_{i}}$ if $i\neq j$ and $x_{j}:=(t-r_{j,b})/r_{j,a_{j}}$. Then $x_{1},\dots,x_{\ell}$ are non-negative integers and $t=x_{i}r_{i,a_{i}}$ if $i\neq j$ and $t=x_{j}r_{j,a_{j}}+r_{j,b}$. In other words, the condition $(ii)$ of Proposition 8.3 is true. Then any projective $\Lambda$-module is a direct sum of finitely generated modules. $(ii)$ then follows from Theorem 4.11.
Now, assume that $(ii)$ is true, and suppose that $\operatorname{Tr}_{R}(M)=R$, and $R$ has Krull dimension $1$. Then, by Lemma 9.2, $\Lambda$ is locally semiperfect. Keeping the notation of Proposition 8.3, we have to prove that $r_{i,a}$ and $r_{j,b}$ are coprime whenever $i\neq j$, $a\in\{1,\dots,t_{i}\}$, $b\in\{1,\dots,t_{j}\}$. Note that it is trivial if either $r_{i,a}=1$ or $r_{j,b}=1$. On the other hand, if $r_{i,a}>1$ then there exists $c\in\{1,\dots,t_{i}\}$ such that $r_{i,c}=1$. Then $r_{i,c}+xr_{i,a}=yr_{j,b}$ can be solvable only if $r_{i,a},r_{j,b}$ are coprime. Now $(i)$ follows from Proposition 8.3.